### v2

# Price Experimentation and Interference

[…]
Price Experimentation and Interference
[…]
In this paper, we examine biases arising in A/B tests where firms modify a continuous parameter, such as price, to estimate the global treatment effect of a given performance metric, such as profit. These biases emerge in canonical experimental estimators due to interference among market participants. We employ structural modeling and differential calculus to derive intuitive characterizations of these biases. We then specialize our general model to a standard revenue management pricing problem. This setting highlights a key pitfall in the use of A/B pricing experiments to guide profit maximization: notably, the canonical estimator for the expected change in profits can have the wrong sign. In other words, following the guidance of canonical estimators may lead firms to move prices in the wrong direction, inadvertently decreasing profits relative to the status quo. We apply these results to a two-sided market model and show how this “change of sign” regime depends on model parameters such as market imbalance, as well as the price markup.Finally, we discuss structural and practical implications for platform operators.
[…]
1 Introduction
In this work we analyze the biases that arise when a technological firm or platform conducts A/B tests by varying a continuous parameter to determine the global treatment effect associated to a performance metric of interest. These biases appear when using canonical designs and estimators and result from interference between market participants.
More concretely, consider the example of a canonical pricing experiment where a fraction of the products are assigned a price that is higher than the baseline price. The global treatment effect of this higher price compares two counterfactual universes: one where all products have the high price (global treatment), and one where all products have the low price (global control). Observe that in the experiment, some consumers who would have purchased at the high price in global treatment instead purchase control units at the low price; when using canonical estimators (such as the difference-in-means estimator), this substitution within the experiment biases the estimation of the global treatment effect. This phenomenon is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts; see Related Work (Section 2) for background.
Our paper considers the following question: Given the standard experimental designs (e.g., product-level or customer-level randomization) and estimators (i.e., difference-in-means) that are employed by platforms, what are the consequences of interference bias in pricing experiments? There are several key features of our approach. First, drawing inspiration from recent work in the operations literature, we employ structural modeling to gain valuable insights into biases. Furthermore, our analysis focuses on differential changes in continuous variables such as prices, enabling us to leverage calculus to derive straightforward and intuitive expressions. (In practical settings, we will think of “small” changes around the control value.) Finally, unlike many preceding studies on online platform experimentation, we focus on profits as an outcome metric, in addition to bookings or transaction volume.
We first introduce a succinct yet general model of platform experimentation in which a firm wants to understand the effect of a continuous variable on a metric of interest. The model is general in that can it encode different market arrangements and experimental designs. By leveraging differential calculus, we provide several structural results regarding the performance of canonical estimators and designs. Notably, we show that the bias corresponds to the cross-effects of the control and treatment metrics with respect to marginal changes on the other group’s variable. These cross-effects are precisely an indicator of interference, providing an intuitive structural characterization of bias.
Second, we specialize our results to a canonical revenue management pricing problem. By using our bias characterizations, we show that experiments can obtain the wrong sign when the metric of interest is profit (or revenue). In particular, we show that when treatment involves raising prices (resp., lowering prices), the global treatment effect may be positive (resp., negative), but due to interference, the estimator can be negative (resp., positive). (Interestingly, when treatment involves raising prices (resp., lowering prices), and the global treatment effect is negative (resp., positive), the experimental estimate obtains the correct sign.) We express the conditions under which an estimate with incorrect sign can be obtained with straightforward expressions that involve price elasticities. [^1] The expressions are reminiscent of elasticity conditions in Blake and Coey, 2014a . Our results have important implications for price optimization based on experiments: incorrectly following the experiment’s suggestion may lead one to move in the wrong direction and decrease profits relative to the status-quo!
We apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022). We show how the biases of listing and customer-side experiments depend on model parameters with a particular focus on regimes in which this “change of sign” occurs. For example, for a listing-side experiment a platform is more likely to obtain the incorrect sign in an experiment if the demand is relatively low. We also calibrate this stylized model to experimental estimates from Airbnb (published in Holtz et al., 2020), and investigate empirical implications for regimes where a change of sign may occur in a price experiment.
2 Related Work
The phenomenon we study is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts. More generally, this interference effect is a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Imbens and Rubin, 2015). Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014b , Fradkin, 2019, Liu et al., 2020), showing its importance in practice. Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, farias2023correcting), and message passing (shirani2023causal).
A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), roemheld2023interference, Deng and Le, (2023). Both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects). roemheld2023interference uses an empirical study to suggest false positives may be inflated due to interference in pricing experiments. More similar to our findings, Deng and Le, (2023) also notes the possibility that the sign of the treatment effect can flip; in their setting this is due to an effect they call lead-day bias in markets with expiring goods. In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign. Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper – those where a continuous variable (such as price) is incrementally varied – as marginal policy changes, and develops a framework to study them.
The closest of these works to ours is Munro et al., (2021), which studies general estimands in an economic price equilibrium setting, and decomposes the treatment effect into direct and indirect effects; the former is obtained via “naive” estimators, while the latter corresponds to our bias, and the authors show how it can be corrected using local price experiments. In that work, the direct and indirect effects are obtained as marginal changes with respect to the treatment assignment probability (inspired by work in Hu et al., (2022)), rather than the price in our present paper. Further, our emphasis in applying our results is on structural insight regarding when changes of sign may be obtained for objectives such as profits or revenue, while their paper is more focused on design and estimation for general estimands in market equilibrium. (An earlier version of Munro et al., (2021) considers treatment changes in a continuous variable, with an emphasis on relating the resulting differential treatment effect to direct and indirect effects.)
3 General Model Formulation
This section presents our general model, which we subsequently adapt to specific settings.
3.1 Market Model and Estimand
We consider a market in which a firm or platform sets a single variable of interest, $x\in\mathbb{R}_{+}$, evaluated by a performance metric $T(x)\in\mathbb{R},\forall x\in\mathbb{R}_{+}$. The performance metric can describe bookings, profits, or revenues of the platform. We make the following assumption that is kept throughout the paper. {assumption} The function $T(x)$ is differentiable for all $x\in\mathbb{R}_{+}$.
The firm is interested in optimizing the performance metric by adjusting the decision variable $x$. In particular, the firm is interested in the difference in the metric when the decision variable takes on a new “treatment” value, denoted by $x_{1}$, relative to a status-quo “control” value, denoted by $x_{0}$. This difference, $T(x_{1})-T(x_{0})$, corresponds to the global treatment effect ($GTE$). We define a local version of $GTE$ considering that the treatment value of the decision variable is “close” to the control value:
$GTE(x_{0})$ ::=:= $T^{\prime}(x_{0})\ .$ (1)
$GTE(x_{0})$ is our estimand.
3.2 Experiment: Design, Model, and Estimator
Our work focuses on settings where the function $T$ may be unknown to the firm or platform. Therefore, an experiment is conducted to determine the direction in which the decision variable should be modified to optimize the performance metric. We assume that the platform conducts an experiment with a single treatment condition where a fraction $q$ of the units are randomized to treatment and a fraction $1-q$ to control. We carry out our analysis in a formal “mean field” or continuum regime, where a continuum mass of units are allocated in each group; this allows us to focus on the magnitude and sign of bias (see also Johari et al., (2022), Wager and Xu, (2021) for related works using a similar approach).
It is worth noting that in the settings we are considering, randomization can occur at either side of the market, i.e., customers or products (i.e., listings in the case of a platform) can be randomized. In a customer-side experiment, a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition. On the other hand, in a product (listing)-side experiment, a customer is exposed to a mix of treatment and control products (or listings). As before, we denote by $x_{0}$ the baseline (control) value of the decision variable and by $x_{1}$ the treatment value.
We assume that the performance metric under the treatment and control conditions are given by the functions $T_{0}(x_{0},x_{1},q)$ and $T_{1}(x_{0},x_{1},q)$, which we refer to as the control and treatment performance metrics, respectively. The function $T_{0}(x_{0},x_{1},q)$ (resp., $T_{1}(x_{0},x_{1},q)$) corresponds to the performance metric in the control (resp., treatment) group when the control and treatment value are given by $x_{0}$ and $x_{1}$, respectively, and a fraction $q$ of units is randomized into treatment. The nature of these functions will become more clear in the context of the applications described below. We make the following assumption that is kept throughout the paper: {assumption} The functions $T_{0}(x_{0},x_{1},q)$, and $T_{1}(x_{0},x_{1},q)$ are differentiable for all $x_{0},x_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
The canonical experimental designs at each side of the market have an associated natural “naive” local estimator. To define this estimator, consider a sequence of experiments for which the treatment value of the decision variable approaches the control value. Then, we compute the relative difference between the performance metrics appropriately normalized by the treatment and control fractions, respectively, which corresponds to the Horvitz-Thompson estimator, or equivalently in our formal mean-field setting, the difference-in-means estimator. More formally, we consider a sequence of experiments indexed by $n\in\mathbb{Z}_{+}$, for which the treatment approaches control: $\lim_{n\to\infty}x_{1}^{n}=x_{0}$. We define a local version of the naive estimator as:
$\widehat{GTE}(x_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$ (2)
We make the following assumption that is kept throughout the paper:
{assumption}
$T_{0}(x,x,q)/(1-q)=T_{1}(x,x,q)/q=T(x)$, for all $x\in\mathbb{R}_{+}$ and $10<q<$ 1.
The assumption states that if the values of control and treatment decision variables are the same, then both the scaled control and treatment performance metrics correspond to the market-wide metric. The assumption will play a key role in our analysis. To further understand it, consider $T_{1}$ with an extended domain $0<q\leq 1$ (the argument for $T_{0}$ is symmetric). Now, suppose $T_{1}$ is linear in its third argument over the relevant domain, that is, for any $0<q\leq 1$ and $10<\alpha<$ 1, $T_{1}(x,x,\alpha q)=\alpha T_{1}(x,x,q)$. If we take $q=1$, then $T_{1}(x,x,\alpha)/\alpha=T_{1}(x,x,1)=T(x)$, where the second equation follows directly by definition. Hence, the assumption is implied by constant (i.e., linear) returns of the performance metric in the fraction randomized. Note that experimental estimates in practice typically implicitly rely on the assumption of constant returns to extrapolate estimates for control and treatment groups to the entire market. For the marketplace models and metrics we study later, this assumption will be satisfied.
Now, we obtain:
$\widehat{GTE}(x_{0},q)$ $=$ $\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{1}(x_{0},x_{0},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$ (3)
$=$ $\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{0}(x_{0},x_{0},q)/(1-q)-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$
$=$ $T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q),$
The first equation is direct from the definition of $\widehat{GTE}(x_{0},q)$, the second follows by Assumption 3.2, and the third by Assumption 3.2. The subscript $x$ (resp., $y$) denotes the partial derivative with respect to the first (resp., second) argument of the function.
Finally, we define the estimation bias of this local estimator as:
$Bias(x_{0},q)$ ::=:= $GTE(x_{0})-\widehat{GTE}(x_{0},q)$ (4)
$=$ $T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)$ (5)
Studying bias in this local regime proves fruitful, as it leads to the following simple characterization.
Proposition 3.1
The bias is given by the following expression:
$Bias(x_{0},q)=T_{0y}(x_{0},x_{0},q)/(1-q)+T_{1x}(x_{0},x_{0},q)/q$
Proof 3.2
Proof. By Assumption 3.2, we have that:
$T(x_{0})=T_{1}(x_{0},x_{0},q)/q\ .$
By Assumption 3.2 and application of the chain rule we get:
$T^{\prime}(x_{0})=T_{1x}(x_{0},x_{0},q)/q+T_{1y}(x_{0},x_{0},q)/q\ .$
Replacing in (5), we obtain the desired result:
$Bias(x_{0},q)=T_{0y}(x_{0},x_{0},q)/(1-q)+T_{1x}(x_{0},x_{0},q)/q\ .$
The result provides an intuitive characterization of bias. Specifically, $T_{0y}(x_{0},x_{0},q)$ and $T_{1x}(x_{0},x_{0},q)$ correspond to the relative changes in performance for the control and treatment groups, respectively, when the decision variable on the other group changes marginally. The cross-partial derivatives represent the effect on the performance metric across experimental groups, making them an intuitive indicator of interference. As interference increases, bias correspondingly intensifies.
4 Revenue Management Setting
We examine a traditional revenue management scenario involving monopoly pricing. A firm or platform offers products that, at least based on observable characteristics, are deemed homogeneous and are thus priced uniformly. The firm faces an unknown to them demand function given by $D(p)$ as a function of the price $p$. The firm has a constant marginal production cost, $c\geq 0$. We make the following assumption that is kept throughout the paper and will imply Assumption 3.1 for the performance metrics we study.
{assumption}
The function $D(p)$ is differentiable and $D^{\prime}(p)<0$, for all $p\in\mathbb{R}_{+}$. At this level of abstraction, the demand function can describe a scenario with either an infinite supply or a limited supply, where a rationing mechanism determines the consumption. In a large limit model, even with supply constraints we expect the demand function to be differentiable.
To provide a concrete example with infinite supply, the demand function can be understood as the fraction among a large population of customers buying the product at a given price. That is, $D(p)=\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}x_{i}(p)$ almost surely, where $\{x_{i}(p):i=1,...,n\}$ are i.i.d. random variables equal to one if customer $i$ buys the product at a given price $p$, and zero otherwise.
4.1 Estimands
We will consider two performance metrics of typical interest: (1) demand itself, which serves as a measure of growth, and (2) the firm’s profits (or revenue, in case costs are zero). The firm’s decision variable is price, and it is seeking to gain an understanding of how changes in price impact demand (i.e., price elasticity of demand) and subsequently affect profit. To achieve this goal, the firm is interested in estimating the difference in demand or profits generated when products are priced at a new “treatment” price relative to the demand or profits generated when products are priced at the status-quo “control” price. These differences correspond to the global treatment effects of interest that we derive below. To do so, we first define the “global” profit function as:
$\pi(p):=(p-c)D(p)\ .$
Following equation (3.1), we present our estimands. These are the local versions of the global treatment effects associated to demand and profits:
$GTE_{D}(p)$ ::=:= $D^{\prime}(p)$ (6)
$GTE_{\pi}(p)$ ::=:= $\pi^{\prime}(p)$ (7)
$=$ $D(p)+(p-c)D^{\prime}(p)$
Note that $GTE_{\pi}(p)$ is the marginal profit of the firm at price $p$.
4.2 Experiment: Design, Model, and Estimator
Given that the demand function is unknown, we assume the firm uses price experimentation to estimate the $GTE$s of interest. As mentioned above, canonical A/B-tests can randomize at either side of the market, and our model does not preclude either possibility. That said, we do note that when conducting price experiments in practice, companies prefer to refrain from randomizing prices on the customer side to adhere to the basic principle that each customer should be offered the same price for the same product (Phillips, 2021).
We assume a single treatment condition in which a fraction $q$ of the subjects are randomized to treatment and a fraction $1-q$ to control. We denote by $p_{0}$ the baseline (control) price and by $p_{1}$ the treatment price. We assume that the experimental demands of control and treated subjects are given by the functions $Q_{0}(p_{0},p_{1},q)$ and $Q_{1}(p_{0},p_{1},q)$, respectively. We further define the functions:
$D_{0}(p_{0},p_{1},q)=\frac{Q_{0}(p_{0},p_{1},q)}{1-q};\quad D_{1}(p_{0},p_{1},q)=\frac{Q_{1}(p_{0},p_{1},q)}{q},$
which correspond to the demand for control and treated subjects scaled by the control and treatment fractions, respectively. Henceforth, we refer to $D_{0}$ and $D_{1}$ as the control and treatment demand functions. Again, at this level of abstraction these demand functions could, for example, incorporate supply constraints and rationing mechanisms.
Recall that in a customer-side experiment treatment and control customers will observe all products priced at the treatment and control prices, respectively, with the proportion of customers assigned to the treatment group being $q$. In contrast, in a product-side experiment, a customer will face an array of products, a fraction $q$ of which are priced at $p_{1}$ and the rest at $p_{0}$.
We make the following assumption that implies Assumption 3.2 and that is kept throughout the paper. {assumption} The functions $D_{0}(p_{0},p_{1},q)$, and $D_{1}(p_{0},p_{1},q)$ are differentiable for all $p_{0},p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1. In addition, $p_{0},p_{1}\geq c$. Note that the assumption precludes the case of perfect substitution between control and treatment products. Under perfect substitution, a negligible undercut of the treatment relative to the control price would result in a discrete increase of the treatment demand; as a result, the demand function would not be continuous. Even though, as initially stated, the firm is selling a single type of product, we can rationalize the smoothness assumption in at least one of two ways: (1) products can be identical in terms of observable characteristics but there may be latent horizontal differentiation; (2) there could be search frictions that do not allow customers to always buy the lowest-priced product.
We introduce the following additional assumption on experimental demand functions that implies Assumption 3.2. {assumption} The following holds:
  1. 1.
    $D_{0}(p_{0},p_{1},q)$ is non-decreasing in $p_{1}$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
  2. 2.
    $D_{1}(p_{0},p_{1},q)$ is non-decreasing in $p_{0}$, for all $p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
  3. 3.
    $D_{0}(p,p,q)=D_{1}(p,p,q)=D(p)$, for all $p\in\mathbb{R}_{+}$ and $10<q<$ 1.
The first two assumptions are natural as they imply that control and treatment products are (imperfect) substitutes. The third assumption states that if the control and treatment prices are the same, then both the scaled control and treatment demand functions are equivalent to the market demand function; this assumption is implied by constant returns to scale. This assumption is satisfied by demand systems for which the control and treatment demand quantities are proportional to their shares in the marketplace. Indeed, this assumption will be satisfied for the model we later study in Section 6, as well for a number of other common demand systems.
We finish this subsection by defining the experimental estimators. To do so, we define the experimental control and treatment profit functions, respectively:
$\pi_{0}(p_{0},p_{1},q)$ ::=:= $(p_{0}-c)D_{0}(p_{0},p_{1},q)$
$\pi_{1}(p_{0},p_{1},q)$ ::=:= $(p_{1}-c)D_{1}(p_{0},p_{1},q).$
Following definition (2) we define naive estimators for demand and profits associated to the canonical designs:
$\widehat{GTE}_{D}(p_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{D_{1}(p_{0},p_{1}^{n},q)-D_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}$
$\widehat{GTE}_{\pi}(p_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{\pi_{1}(p_{0},p_{1}^{n},q)-\pi_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}$
Note that analogs of these naive estimators are frequently used in practice in estimating global treatment effects from experimental data.
Now, by equation (3), we obtain:
$\widehat{GTE}_{D}(p_{0},q)$ $=$ $D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\ ,$ (8)
$\widehat{GTE}_{\pi}(p_{0},q)$ $=$ $\pi_{1y}(p_{0},p_{0},q)-\pi_{0y}(p_{0},p_{0},q)$ (9)
$=$ $D(p_{0})+(p_{0}-c)\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\ ,$
where we used Assumption 4.2 and the last equation follows by Assumption 4.2.3.
5 Biases: Analysis and Results
In this section we analyze biases associated to the naive estimators. Following definition (4) we define the estimation bias for demand and profits associated to the canonical designs:
$Bias_{D}(p_{0},q)$ ::=:= $GTE_{D}(p_{0})-\widehat{GTE}_{D}(p_{0},q)$
$Bias_{\pi}(p_{0},q)$ ::=:= $GTE_{\pi}(p_{0})-\widehat{GTE}_{\pi}(p_{0},q)$
First, note that by Proposition 3.1 we obtain the following bias characterizations:
$Bias_{D}(p_{0},q)$ $=$ $D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)$ (10)
$Bias_{\pi}(p_{0},q)$ $=$ $\pi_{0y}(p_{0},p_{0},q)+\pi_{1x}(p_{0},p_{0},q)$ (11)
$=$ $(p_{0}-c)\left(D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\right)$
$=$ $(p_{0}-c)Bias_{D}(p_{0},q)$
As discussed in Proposition 3.1, bias increases with increasing cross-effects between treatment and control groups. More specifically, the bias is influenced by the degree to which the control (or treatment) group detracts demand from its counterpart in response to a marginal price change in the latter group.
Now, note that under Assumptions 4.2 and 4.2, both $Bias_{D}(p_{0},q)$ and $Bias_{\pi}(p_{0},q)$ are non-negative, so we obtain that the naive estimators always (weakly) underestimate $GTE$:
Proposition 5.1
We have that $\widehat{GTE}_{D}(p_{0},q)\leq GTE_{D}(p_{0})$ and $\widehat{GTE}_{\pi}(p_{0},q)\leq GTE_{\pi}(p_{0})$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
The result follows similar cannibalization arguments developed in previous work (Johari et al., 2022). To provide intuition consider a listing-side experiment in which treatment price is larger than control price. In this case, experimental treatment demand underestimates global treatment demand because of substitution in the experiment from the treatment to the control group. Similarly, experimental control demand overestimates global control demand. Now, let’s consider a customer-side experiment where there are supply constraints, so customers compete for limited products. In such cases, the control customers may “cannibalize” supply from the treatment group, given the lower control price, leading again to an underestimation of the global treatment effect. Note that in a customer-side experiment without supply constraints, we anticipate no bias because there is no competition between the treatment and control groups.
It is important to observe that because $D^{\prime}(p)<0$ by Assumption 4, $GTE_{D}(p_{0})<0$, for all $p$. Moreover, since $\widehat{GTE}_{D}(p_{0},q)$ is an underestimation, both the estimand and the naive estimator will have the same (nonpositive) sign.
However, there is a more intriguing result for profits. Suppose that the price $p_{1}$ is greater than $p_{0}$. In this scenario, even if the demand $D(p_{1})$ at $p_{1}$ is lower than the demand $D(p_{0})$ at $p_{0}$, the increase in price may offset the decrease in demand, leading to a positive true global treatment effect in profits. However, because the naive estimator underestimates the treatment effect, it could be negative. Next, we indeed provide conditions under which the naive estimator not only underestimates the global treatment effect, but also obtains the wrong sign.
By definition, $GTE_{\pi}(p_{0})\geq 0$ if and only if $\pi^{\prime}(p_{0})\geq 0$, which after some simple manipulation yields the following standard “markup” inequality:
[…]
where $e_{p}(p_{0})=D^{\prime}(p_{0})p_{0}/D(p_{0})$ is the price elasticity of demand. If $\pi(p)$ is strictly quasi-concave, the condition is equivalent to the price $p_{0}$ being below the optimal monopoly price. For the case $c=0$, condition (a) naturally states that marginal revenue is increasing at $p_{0}$, or equivalently that the demand is relatively inelastic with respect to price.
Now, consider the naive estimator. Note that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ is equivalent to $Bias_{\pi}(p_{0},q)\geq{GTE}_{\pi}(p_{0})$. That is, provided that condition $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ holds, we have a change of sign when the bias is so severe that it becomes larger than the treatment effect. Furthermore, after some manipulation of equation (9), we obtain that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ if and only if:
[…]
Note that condition (b) is a “modified markup” inequality. We have the following result:
Proposition 5.2
  1. 1.
    Condition (a) holds if and only if $GTE_{\pi}(p_{0})\geq 0$.
  2. 2.
    Condition (b) holds if and only if $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$.
  3. 3.
    $A(p_{0})\leq B(p_{0},q)$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
Proof 5.3
Proof. Points (1) and (2) in the proposition are direct from the arguments above. To prove point (3), note that $D_{1y}(p_{0},p_{0},q)\leq D^{\prime}(p_{0})$ (by applying Assumptions 4.2.2 and 4.2.3 together with the chain rule) and $D_{0y}(p_{0},p_{0},q)\geq 0$ (by Assumption 4.2.1). Hence,
$\frac{p_{0}-c}{p_{0}}\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\leq\frac{p_{0}-c}{p_{0}}D^{\prime}(p_{0})\ ,$
and thus the result follows because $D^{\prime}(p_{0})<0$ by Assumption 4.
The proposition states a sign change is obtained, $\widehat{GTE}_{\pi}(p_{0},q)\leq 0\leq GTE_{\pi}(p_{0})$, if and only if conditions (a) and (b) hold, that is there are appropriate lower and upper bounds on the inverse of the price elasticity. Note that we use weak inequalities in conditions (a) and (b) for a change of sign; strictly speaking of course, a change of sign can only occur if at least one of the two inequalities is strict. [^2] If exactly one of these inequalities is strict (and both hold), the canonical estimator will still instruct firms to move in a non-optimal direction; either by telling them to lower prices when they were already at the optimal price, or by suggesting they keep their price constant when they were underpricing. We suppress this edge case in the sequel in favor of simplicity of presentation. [^3] In our two-sided market model, introduced in section 6, the scenario $\widehat{GTE}_{\pi}(p_{0},q)=GTE_{\pi}(p_{0})=0$ is impossible, meaning this edge case does not exist.
To provide intuition, note that by using equation (8) we can define the experimental price elasticity of demand as:
$\widehat{e_{p}}(p_{0},q)=\frac{p_{0}(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q))}{D(p_{0})}.$ (12)
Hence, we can re-write condition (b) as:
$-\frac{1}{\widehat{e_{p}}(p_{0},q)}\leq\frac{(p_{0}-c)}{p_{0}}\ ,$
which can be understood as the experimental analogue of a markup inequality. It states that at the experimental elasticity, the current price is above the optimal monopoly price (under an appropriate concavity assumption). Hence, as expected, a change of sign occurs when the naive estimator indicates that the platform is overpricing relative to the optimal monopoly price (and therefore prefers to decrease prices), when in reality it is underpricing (and therefore should raise prices). The result has important consequences for experiment-driven price optimization: the prescription implied by the experiment (reduce price) is the opposite to the one implied by the $GTE$ (increase price). Hence, if one moves in the direction suggested by the experiment, profits decrease relative to the status-quo.
To provide further intuition, let us assume that condition (a) holds and that profits are quasiconcave, so that the firm is underpricing relative to the optimal monopoly price. Hence, we have a change of sign if and only if condition (b) holds. If $|D_{1y}(p_{0},p_{0},q)|+D_{0y}(p_{0},p_{0},q)$ is large compared to $|D^{\prime}(p_{0})|$, condition (b) will hold. Note that $|D_{1y}(p_{0},p_{0},q)|$ represents the magnitude of the decrease in treatment demand given a small increase in treatment price, which captures two effects: (i) demand lost to the outside option; and (ii) cannibalization to the control group. Note that (i) is the global demand effect we aim to estimate. On the other hand, the term $D_{0y}(p_{0},p_{0},q)$ represents the increase in control demand given a small increase in treatment price, which captures cannibalization from the treatment to the control group (equivalent to effect (ii)). A change of sign happens as the two cannibalization effects just mentioned become larger.
We conclude with a remark regarding our estimators, our findings on changes of sign, and guidance for practitioners. Note that the estimators $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that we defined are estimates of derivatives of the demand and profit function, respectively, as is natural given our estimands. In practice, a platform will typically directly use a difference in estimated treatment and control demand, or estimated treatment and control profit, without normalizing by the difference in prices. This can have some bearing in interpreting our results, depending on the sign of the difference between the treatment and control price.
Concretely, suppose that treatment involves raising the price relative to control. Then our results show that a change of sign can only occur if the true global treatment effect on profits is positive, but the treatment effect is estimated to be negative in the experiment. In other words, in this case, the platform risks failing to move prices in a profitable direction.
On the other hand, suppose the treatment involves lowering the price relative to control. Then our results show that a change of sign can only occur if the true global treatment effect on profits is negative, but the treatment effect is estimated to be positive in the experiment. In other words, in this case, the platform risks moving prices in a potentially unprofitable direction.
6 A Two-Sided Market Model
In this section, we illustrate the application of our results of the general revenue management setting introduced in Section 4 to a two-sided market model developed in Johari et al., (2022). That paper develops a stochastic dynamic market model with supply constraints in which customers book listings for a finite period of time. The paper also introduces a mean-field limit that greatly simplifies the analysis. We focus on the mean-field model here and describe its main elements.
6.1 Preliminaries
We begin by recapping the mean-field model formulation in Johari et al., (2022), adapted to our setting with price experiments. There is a two-sided platform evolving in continuous time with listings on the supply-side and customers on the demand side. There is a continuum of listings of mass $\rho$ that are homogenous with respect to their observable characteristics. Having a limited supply will play a key role in the analysis. We assume all listings have the same price $p$. At each time $t$, listings can be available or booked. The state of the system at time $t$ is given by $s_{t}$, which denotes the mass of available listings available at time $t$.
[…]
$q(s)=\frac{sv(p)}{\varepsilon+sv(p)}.$
Finally, a listing remains occupied for an expected time $\frac{1}{\tau}$ when it gets booked; in our analysis we normalize $\tau$ to be 1 without loss of generality. Note that in our model, the parameter $\lambda$ mediates market balance: when $\lambda\rightarrow 0$ the market becomes demand constrained, when $\lambda\rightarrow\infty$ the market becomes supply constrained.
Note that the general model in Johari et al., (2022) allows for heterogeneity in both listings and consumers. In the model we consider in this section, we make both consumers and listings homogeneous.
As in Johari et al., (2022) we focus on the steady-state of the system. Let $s^{*}(\lambda,p)$ be the steady-state mass of available listings for a given market balance parameter $\lambda$ and price $p$. It is shown in Johari et al., (2022) that $s^{*}(\lambda,p)$ is the unique solution to the following conservation of flow equation:
$\rho-s^{*}(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)v(p)}{\varepsilon+s^{*}(\lambda,p)v(p)}:=D(\lambda,p)\ .$ (13)
The left-hand side of the equation corresponds to the rate at which listings become available. The right-hand side corresponds to the booking (or demand) rate, defined as $D(\lambda,p)$. In the steady-state both rates are equal.
{assumption}
We assume the following for the utility function $v(p)$:
  1. 1.
    $v(p)$ is strictly positive for $p\in[c,+\infty);$
  2. 2.
    $v(p)$ is strictly decreasing for $p\in[c,+\infty);$
  3. 3.
    $v(p)$ is differentiable for $p\in[c,+\infty)$, and $v^{\prime}(p)$ is continuous;
  4. 4.
    $-pv^{\prime}(p)/v(p)$ is strictly increasing in $p$ for $p\in[c,+\infty).$
[…]
6.2 Steady-State and Estimand
To gain some initial intuition, we start this section by providing some analysis on the steady-state and the associated demand. Then, we define our estimands.
First, we note that listing availability $s^{*}(\lambda,p)$ is monotonically increasing with price $p$ and decreasing with the market balance parameter $\lambda$, which is natural. These monotonicity results can be derived by differentiating $s^{*}(\lambda,p)$ in equation (13). [^7] Indeed, we obtain $s^{*}_{p}(\lambda,p)=-\frac{s^{*}(\lambda,p)v^{\prime}(p)\varepsilon}{\frac{1}{\lambda}(\varepsilon+s^{*}(\beta,p)v(p))^{2}+v(p)\varepsilon}>0$ and $s^{*}_{\lambda}(\lambda,p)=-\frac{v(p)s(\lambda,p)(v(p)s(\lambda,p)+\varepsilon)}{\varepsilon v(p)(2s(\lambda,p)+\lambda)+v(p)^{2}s(\lambda,p)^{2}+\varepsilon^{2}}<0$. Figure 4 in the Appendix illustrates the steady-state listing availability as a function of price ($p$) and market balance ($\lambda$).
Naturally, $D(\lambda,p)$ is monotonically decreasing in price ($p$) and increasing in market balance ($\lambda$). Again, we can prove these monotonicity results by differentiating. [^8] This can be seen by rewriting the demand as the rate at which the listings become available. Using the market balance equation, we have $D(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)v(p)}{\varepsilon+s^{*}(\lambda,p)v(p)}=\rho-s^{*}(\lambda,p)$. Hence, $D_{p}(\lambda,p)=-s^{*}_{p}(\lambda,p)$ and $D_{\lambda}(\lambda,p)=-s^{*}_{\lambda}(\lambda,p)$. Since $s^{*}_{p}(\lambda,p)>0$ and $s^{*}_{\lambda}(\lambda,p)<0$ we obtain that $D_{p}(\lambda,p)<0$ and $D_{\lambda}(\lambda,p)>0$. Figure 5 in the Appendix shows the dependence of the demand function on the market balance parameter and price.
Now, we define our estimands. Following the definitions in equations (6) and (7):
$GTE_{D}(\lambda,p)$ $=D_{p}(\lambda,p)$
$GTE_{\pi}(\lambda,p)$ $=D(\lambda,p)+(p-c)D_{p}(\lambda,p)$
$=\rho-s^{*}(\lambda,p)-(p-c)s^{*}_{p}(\lambda,p),$ (14)
where in the last equation we used equation (13).
6.3 Experimental Designs
In this section we analyze two canonical experimental designs in platforms: randomizing listings or randomizing customers.
6.3.1 Listing-Side Randomization
We assume that randomization occurs on the listing side, where a fraction $q$ of the listings is assigned to treatment. Hence, a customer is exposed to a mix of treatment and control listings. In this experimental design, the balance equations are given by:
$(1-q)\rho-s_{0}^{*}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s_{0}^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s_{0}^{*}(\lambda,p_{0},p_{1},q)v(p_{0})+s_{1}^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}$ (15)
$q\rho-s_{1}^{*}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s_{1}^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s_{0}^{*}(\lambda,p_{0},p_{1},q)v(p_{0})+s_{1}^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}\ ,$
where $s^{*}_{0}(\lambda,p_{0},p_{1},q)$ and $s^{*}_{1}(\lambda,p_{0},p_{1},q)$ refer to the steady-state available control and treated listings, respectively.
The control and treatment demand functions $D^{LR}_{0}(\lambda,p_{0},p_{1},q)$ and $D_{1}^{LR}(\lambda,p_{0},p_{1},q)$ are given by:
$D_{0}^{LR}(\lambda,p_{0},p_{1},q)$ $=\rho-s_{0}^{*}(\lambda,p_{0},p_{1},q)/(1-q);$ (16)
$D^{LR}_{1}(\lambda,p_{0},p_{1},q)$ $=\rho-s_{1}^{*}(\lambda,p_{0},p_{1},q)/q.$
We can show that $D^{LR}_{0}(\lambda,p_{0},p_{1},q)$ and $D^{LR}_{1}(\lambda,p_{0},p_{1},q)$ satisfy Assumption 4.2 (see Appendix 10.4).
Following the characterization of bias from equations (10) and (11), we obtain that the bias in this setting, $Bias_{\pi}^{LR}(\lambda,p_{0},q)$ , is given by:
$Bias^{LR}_{\pi}(\lambda,p_{0},q)$ $=(p_{0}-c)(D^{LR}_{1x}(\lambda,p_{0},p_{0},q)+D^{LR}_{0y}(\lambda,p_{0},p_{0},q))$
$=(p_{0}-c)\frac{-\lambda s^{*}(\lambda,p_{0})^{2}v(p_{0})v^{\prime}(p_{0})(\varepsilon+s^{*}(\lambda,p_{0})v(p_{0}))}{[\varepsilon+s^{*}(\lambda,p_{0})v(p_{0})+\lambda v(p_{0})][(s^{*}(\lambda,p_{0})v(p_{0})+\varepsilon)^{2}+\lambda v(p_{0})\varepsilon]}$ (17)
Note that the bias is independent of the treatment fraction $q$.
6.3.2 Customer-Side Randomization
In this case we randomize on the customer side, with a fraction $q$ of customers randomized into treatment. In this case, a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition. The steady-state is given by:
$\rho-s^{*}(\lambda,p_{0},p_{1},q)=\lambda\left[q\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}+(1-q)\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}\right].$ (18)
In this setting we have the following demand functions:
$D^{CR}_{0}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})},$ (19)
$D^{CR}_{1}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}.$
One can show that the demand functions $D^{CR}_{0}(\lambda,p_{0},p_{1},q)$ and $D^{CR}_{1}(\lambda,p_{0},p_{1},q)$ satisfy Assumption 4.2 (see Appendix 10.5).
Using the characterization in equation (11) we obtain the following expression for $Bias_{\pi}^{CR}$:
$Bias^{CR}_{\pi}(\lambda,p_{0},q)$ $=(p_{0}-c)(D^{CR}_{1x}(\lambda,p_{0},p_{0},q)+D^{CR}_{0y}(\lambda,p_{0},p_{0},q))$
$=-(p_{0}-c)\frac{\lambda^{2}\varepsilon^{2}v(p_{0})v^{\prime}(p_{0})s^{*}(\lambda,p_{0})}{(v(p_{0})s^{*}(\lambda,p_{0})+\varepsilon)^{2}\left[(s^{*}(\lambda,p_{0})v(p_{0})+\varepsilon)^{2}+\lambda v(p_{0})\varepsilon\right]}$ (20)
Note that the bias in the CR-setting is independent of the treatment fraction $q$.
6.4 Properties of Biases with Market Balance
In this section we study structural properties of biases with a particular focus on extreme regimes of market balance. The analysis in this subsection closely follows Johari et al., (2022).
[…]
As the market becomes more demand constrained ($\lambda\rightarrow 0$), competition between customers declines resulting in less interference in a CR experiment. In the limit, there is no competition between customers and we expect the bias in the CR experiment to be zero. On the other hand, competition between listings increases resulting in more interference in a LR experiment. In the limit, we expect the bias in the LR experiment to be non-zero.
Conversely as the market becomes supply constrained ($\lambda\rightarrow\infty$), competition between listings declines resulting in less interference in a LR experiment. In the limit, there is no competition between listings and we expect the bias in the LR experiment to be zero. Competition between customers increases resulting in more interference in a CR experiment. In the limit, we expect the bias in the CR experiment to be non-zero.
The following propositions formalize the intuition about the behavior of the bias in both extremes of market balance.
Proposition 6.1
For all $p\in[c,+\infty)$, we have that:
$\lim_{\lambda\to 0}GTE_{D}(\lambda,p)/\lambda=\frac{\rho v^{\prime}(p)\varepsilon}{(\varepsilon+\rho v(p))^{2}};\ \ \lim_{\lambda\to 0}GTE_{\pi}(\lambda,p)/\lambda=\frac{\rho v(p)}{\varepsilon+\rho v(p)}+\frac{(p-c)\rho v^{\prime}(p)\varepsilon}{(\varepsilon+\rho v(p))^{2}}.$
We have that $\lim_{\lambda\to 0}Bias_{\pi}^{CR}(\lambda,p)/\lambda=(p-c)\lim_{\lambda\to 0}Bias_{D}^{CR}(\lambda,p)/\lambda=0$, while
$\lim_{\lambda\to 0}Bias_{\pi}^{LR}(\lambda,p)/\lambda=(p-c)\lim_{\lambda\to 0}Bias_{D}^{LR}(\lambda,p)/\lambda=-\frac{(p-c)\rho^{2}v(p)v^{\prime}(p)}{(\varepsilon+\rho v(p))^{2}}.$
Proposition 6.2
For all $p\in[c,+\infty)$, we have that: $\lim_{\lambda\to\infty}GTE_{D}(\lambda,p)=0$ and that:
$\lim_{\lambda\to\infty}GTE_{\pi}(\lambda,p)=\rho$. We have $\lim_{\lambda\to\infty}Bias_{\pi}^{LR}(\lambda,q)=(p-c)\lim_{\lambda\to\infty}Bias_{D}^{LR}(\lambda,q)=0$, while
$\lim_{\lambda\to\infty}Bias_{\pi}^{CR}(\lambda,p)=(p-c)\lim_{\lambda\to\infty}Bias_{\pi}^{CR}(\lambda,p)=-\frac{\rho(p-c)v^{\prime}(p)}{v(p)}.$
In the limit where $\lambda\rightarrow 0$, Proposition 6.1 suggests that the CR experiment is unbiased while the LR experiment is biased. Similarly, in the limit where $\lambda\rightarrow\infty$, Proposition 6.2 suggests that the CR experiment is biased while the LR experiment is unbiased. Proofs of both propositions are presented in Appendix 10. The proofs are similar to the analysis in Johari et al., (2022) but we also leverage our characterization of biases based on cross-partial derivatives. In the next section we will further study the extent of these biases to characterize change of sign regimes.
We also provide some figures to complement our limit results. In Figure 1. As $\lambda$ increases, the market becomes more supply-constrained and bias decreases as competition between listings dissipates. Figure 2 shows the normalized bias of profits for CR, and this dependence on $\lambda$ is naturally reversed. Both figures also show interesting behavior of biases with respect to the baseline price $p$. As the price $p$ changes, the sensitivity of demand to price changes as well; in particular, in both CR and LR experiments, both $D_{0y}$ and $D_{1x}$ tend to be smaller at larger prices (i.e., demand becomes more inelastic as the price becomes large); this behavior drives the shape of the bias curve with respect to price, and subsequently profit.
Figure 1: Illustration of $Bias^{LR}_{D}/\lambda$ and $Bias^{LR}_{\pi}/\lambda$ as a function of price ($p$) and market balance ($\lambda$). Instance is given by $v(p)=e^{V-p}$, $\epsilon=1$, $c=1$, $V=5$, and $\rho=1$.
Figure 2: Illustration of $Bias^{CR}_{D}/\lambda$ and $Bias^{LR}_{\pi}/\lambda$ as a function of price ($p$) and market balance ($\lambda$). Instance is given by $v(p)=e^{V-p}$, $\epsilon=1$, $c=1$, $V=5$, and $\rho=1$.
6.5 Change of Sign
We recall the two conditions for a change of sign to occur:
[…]
An interesting immediate observation is that the regimes in which a change in sign occurs in an LR or CR experiment using the corresponding naive estimators are independent of the treatment fraction $q$, because the corresponding biases are independent of $q$. This has important implications for practice, as the treatment fraction is not a lever that can prevent the possibility that the sign of the treatment effect estimate is incorrect.
Using the structure of the model we can provide the following insightful characterizations for the change of sign regions, as follows.
Theorem 6.3
For all $\lambda\in(0,+\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{LR}_{\lambda}\subseteq[c,\infty)$ such that the sign of the listing-side experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p^{\prime}\in[c,+\infty)$ will be wrong if and only if $p^{\prime}\in P^{LR}_{\lambda}$.
Theorem 6.4
For all $\lambda\in(0,+\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{CR}_{\lambda}\subseteq[c,\infty)$ such that the sign of the customer-side experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p^{\prime}\in[c,+\infty)$ will be wrong if and only if $p^{\prime}\in P^{CR}_{\lambda}$.
Theorems 6.3 and 6.4 tell us that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$ such that the sign of the canonical estimator will be wrong, regardless of whether a listing-side or customer-side experiment is run. The proofs are provided in Appendix 10.9 and repeatedly exploit the market balance equations for CR and LR respectively.
The right hand side of both the intervals $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ is the optimal monopoly price $p^{*}$, i.e., where $GTE_{\pi}(\lambda,p^{*})=0$. In particular, this reveals that the sets $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not comprised of unrealistic prices: if a firm is currently pricing optimally or very slightly below optimally, the sign of their canonical estimator will be wrong, regardless of their experimental setup.
It is not true, however, that the two experimental setups are equally fallible in all settings. Note that $\widehat{e_{p}}^{LR}$ is not equal to $\widehat{e_{p}}^{CR}$, meaning that the left endpoints of $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not the same, and behave quite differently based on market balance, as we illustrate through the following numerical study.
6.5.1 Numerical Results
Figure (3) shows numerical results on the set of prices where a change of sign occurs.
Figure 3: Illustration of the change of sign region (white) in an LR-experiment (left) and CR-experiment (right), respectively, with $v(p)=e^{V-p}$, $V=5$, $c=0.5$, $\rho=1$, and $\epsilon=1$. The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in black. The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in grey.
Consistent with Propositions 6.1 and 6.2, we observe that the change of sign region becomes empty in the demand-constrained regime for a CR experiment and in the supply constrained experiment for an LR experiment. At intermediate market balance regimes, we observe that the change of sign region corresponds to a single continuous interval on prices, as the theorems above suggest. We observe that in the demand-constrained regime Condition (b) for a change of sign is more likely to be satisfied by the LR experiments, while in the supply-constrained regime Condition (b) is more likely to be satisfied by the CR experiments as expected. We also see that the region where a change of sign occurs is, overall, increasing with market balance for the CR experiments and decreasing for the LR experiment.
7 An Empirical Case Study
Our results show that price experiments can get the sign wrong, and suggest that practitioners would benefit from investigating whether their experiments are vulnerable to such changes of sign. In this section, we present an empirical illustration to study this question, by calibrating a version of our stylized model to data from Airbnb. Although our model simplifies the complexity of the Airbnb marketplace in many dimensions, we believe this exercise is valuable because it offers practical insights into how our results can align with empirical data. As we will see, our results suggest that a change of sign is indeed plausible in the Airbnb context.
We rely on the estimates by Holtz et al., (2020) of a guest fee “meta-experiment” at Airbnb. This experiment involved a treatment that raised platform fees. The design consisted of two “arms”: randomization across clusters of listings in one arm; as well as Bernoulli randomization across listings in a second arm. Holtz et al., (2020) use this meta-experiment to estimate the degree to which Bernoulli randomized experiments exhibit interference, by comparing with estimates from the cluster randomized arm (which mitigates interference effects). The paper reports on cluster and naive experiment estimates and standard errors for the global treatment effect (the TATE, or total average treatment effect in their paper) for bookings, nights booked, and gross spend.
To translate from Holtz et al., (2020) to our setting, we first recap certain features of their presentation. First, even though the treatment corresponds to a percentage increase in prices, the size of this treatment intervention is not disclosed in Holtz et al., (2020). For our illustrative exercise here, we interpret the estimates of Holtz et al., (2020) as semi-elasticities, i.e., absolute changes in quantities for a 1% price increase. Second, Holtz et al., (2020) does not disclose absolute levels (e.g., of nights booked, or revenue); instead, these quantities are rescaled by the same randomly chosen quantity (not disclosed to the reader). As we will see, this does not present any challenges for our exercise here.
We need to slightly adjust the analysis to accommodate a percentage change in prices relative to the bias analysis in Section 5 that considers an absolute change. Such an analysis can be carried out by considering demand $D(\alpha p)$ and profit $(\alpha p-c)D(\alpha p)$, and considering derivatives with respect to $\alpha$, evaluated at $\alpha=1$. Proceeding in this way, we redefine:
$GTE_{D}(p)=pD^{\prime}(p);\ \ GTE_{\pi}(p)=p(D(p)+pD^{\prime}(p))-cpD^{\prime}(p)$
as the semi-elasticities of demand and profit, respectively.
As before, the first condition for a change of sign at $p_{0}$ is (a) $GTE_{\pi}(p_{0})\geq 0$, which becomes:
$p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))-cp_{0}D^{\prime}(p_{0}).$ (21)
The second condition is that (b) $Bias_{\pi}(p_{0},q)\geq GTE_{\pi}(p_{0})$ (where now we note that the bias is computed using percentage price changes, i.e., with respect to changes in $\alpha$). Since $Bias_{\pi}(p_{0},q)=(p_{0}-c)Bias_{D}(p_{0},q)$, condition (b) becomes equivalent to:
$(p_{0}-c)Bias_{D}(p_{0},q)\geq p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))p_{0}-cp_{0}D^{\prime}(p_{0}).$ (22)
We now leverage the empirical estimates of Holtz et al., (2020). The cluster arm estimates are a decrease of $1.136$ in nights booked per listing, and of $116.67$ in gross revenues. For our analysis, we assume that the cluster arm provides ground truth, and the corresponding bias for demand is reported to be $0.205$. We assume that $D(p_{0})$ represents nights booked per listing and $p_{0}$ the average price per night, so that $p_{0}D(p_{0})$ represents gross revenues per listing. In particular, note that then $p_{0}D^{\prime}(p_{0})$ is the semi-elasticity of nights booked, and $p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))$ is the semi-elasticity of gross revenues per listing. Finally, we assume $c$ represents the marginal (opportunity) cost per night.
By equation (21) and the estimates, the first condition for change of sign amounts to $-116.7+1.136c\geq 0$, because we identify $-116.67$ and $-1.136$ as the ground-truth semi-elasticities of revenues and demand, respectively. This simplifies to
$c\geq 102.73.$ (23)
The second condition for change of sign amounts to $0.205(p_{0}-c)\geq-116.67+1.136c$, which can be simplified to
$116.67+0.205p_{0}\geq 1.341c$ (24)
Importantly, note that all quantities used in these inequalities are re-scaled by the same random constant, and so these inequalities are valid for any value of that constant under our assumptions.
The pair of inequalities (23)-(24) can be used to interrogate whether a change of sign is plausible in this setting. For example, for a reasonable marginal opportunity cost of $c=125$ (corresponding to \$125 per night), our analysis suggests that market segments with a baseline price of $p_{0}=249$ (corresponding to \$249 per night) yield an incorrectly estimated sign. More generally, this exercise allows us to identify combinations of prices and costs for which the canonical estimator would result in a change of sign. In particular, (24) can be rewritten as $0.847-87.002/p_{0}\leq(p_{0}-c)/p_{0}$, lower bounding markups based on observed market prices that yield a change of sign. For example, the average price of recently booked Airbnb listings in San Francisco is given by \$212 (cf http://insideairbnb.com/san-francisco). According to our analysis, this implies a markup of $0.43$ or larger would yield a change of sign. Given that these are markups that are reasonable in practice, we believe this stylized exercise motivates further detailed empirical study of the prevalence of such changes of sign in real-world platforms.
8 Conclusion and Future Work
In this paper we studied biases that arise due to interference in platform experiments, with a particular emphasis on profit or revenue metrics. Notably, we find that in certain settings, these biases can be so severe that they could introduce a “change of sign” in estimation of profit or revenue.
In future work, we aim to extend the analysis to other settings of interest, including two-sided market settings where the variable of interest is the percentage fee charged by the platform (as a fraction of gross revenue). Further, our findings in Section 7 motivate further empirical study of the prevalence of changes of sign in practice. It is natural to consider more robust empirical calibration strategies, as well as practical methodological approaches to mitigating this change of sign. These remain important directions for future work.