### v4

# Price Experimentation and Interference

[…]
Price Experimentation and Interference
[…]
In this paper, we examine the biases that arise when firms run A/B tests on continuous parameters to estimate global treatment effects on performance metrics of interest; we particularly focus on price experiments to measure the price impact on quantity demanded, and on profit. In canonical A/B experimental estimators, biases emerge due to interference between market participants. We employ structural modeling and differential calculus to derive intuitive characterizations of these biases. We then specialize our general model to the standard revenue-management pricing problem. This setting highlights a fundamental risk innate to A/B pricing experiments: that the canonical estimator for the expected change in profits, counterintuitively, can have the wrong sign in expectation. In other words, following the guidance of canonical estimators may lead firms to move prices (or fees) in the wrong direction, inadvertently decreasing profits. We introduce a novel debiasing technique for these canonical experiments, requiring only that firms equally split units between treatment and control. We apply these results to a two-sided market model, and demonstrate how the “change of sign” regime depends on market factors such as the supply/demand imbalance, and the price markup. We conclude by calibrating our two-sided market model to published empirical estimates from Airbnb marketplaces, demonstrating that estimators with the wrong sign are not a knife-edge issue, and that they may be prevalent enough to be of concern to practitioners. [^1] We would like to thank Wassim Dhaouadi for his valuable input during the beginning stages of the project, and Andrey Fradkin for his critical assistance for our Airbnb empirical case study. We are also grateful to Eric Aldrich, Alex Deng, Thu Le, Ruben Lobel, Ali Rauh, Dan Schmierer, and seminar participants at Boston University, University of Chile, the MIT Conference on Digital Experimentation, the INFORMS Revenue Management and Pricing Conferencea, the INFORMS Annual Meeting, the INFORMS Manufacturing and Service Operations Management Conference, and Airbnb for helpful feedback. Orrie Page was supported by a National Science Foundation Graduate Research Fellowship.
[…]
1 Introduction
In this work, we analyze the biases that arise when a firm or platform conducts an A/B test by varying a continuous parameter, in order to determine the global treatment effect associated with a performance metric of interest. These biases appear when using canonical experimental designs and estimators, and are the result of interference between market participants.
More concretely, consider the example of a typical pricing experiment, commonly used by online platforms, where a fraction of the products are assigned a treatment price that is higher than the baseline (control) price. The global treatment effect of this higher price compares two counterfactual universes: one where all products have the high price (global treatment), and one where all products have the low price (global control). Observe that in the experiment, some consumers who would have purchased at the high price in global treatment instead purchase control units at the low price. When using canonical estimators (such as the difference-in-means estimator), this substitution within the experiment biases the estimation of the global treatment effect. This phenomenon is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts; see Related Work (Section 2) for background.
Our paper considers the following question: Given the standard experimental designs (e.g., listing-side randomization, LR, or customer-side randomization, CR) and estimators (i.e., difference-in-means) that are employed by platforms, what are the consequences of interference bias in pricing experiments? There are several key features of our approach. First, drawing inspiration from recent work in the operations literature, we employ structural modeling to gain valuable insights into biases. Furthermore, our analysis focuses on differential changes in continuous variables such as prices, enabling us to leverage calculus to derive straightforward and intuitive expressions. (In practical settings, this corresponds to “small” changes around the control price.) Finally, unlike many preceding studies on online platform experimentation, we focus on profits as an outcome metric, in addition to bookings or transaction volume.
We first introduce in Section 3 a succinct yet general model of platform experimentation in which a firm wants to understand the impact of modifications to a continuous variable, on a metric of interest. The model is general, in the sense that it can represent various market arrangements and experimental designs. By leveraging differential calculus, we provide several structural results regarding the performance of canonical estimators and designs. Notably, we show that the bias corresponds to the cross-effects of the control and treatment group metrics, with respect to marginal changes in the other group’s variable. These cross-effects are precisely an indicator of interference, providing an intuitive structural characterization of bias.
Second, in Sections 4 and 5, we specialize our results to a canonical revenue management pricing problem. By using our bias characterizations, we show that the bias can be severe enough that experimenters can obtain the wrong sign when the metric of interest is profit (or revenue). In particular, we show that when treatment involves raising prices (resp., lowering prices), the global treatment effect may be positive (resp., negative), but due to interference, the estimator can be negative (resp., positive) in expectation. Interestingly, when treatment involves raising prices (resp., lowering prices), and the global treatment effect is negative (resp., positive), the experimental estimate always obtains the correct sign. We express the conditions under which an estimate with the wrong sign can be obtained with straightforward expressions that involve price elasticities. [^2] The expressions are reminiscent of elasticity conditions in Blake and Coey, (2014), though their analysis differs from ours: they focus on estimating the relative bias in terms of supply and demand elasticities, and carry out a local change with respect to a multiplicative factor by which treatment demand changes relative to control. By contrast, our estimand is local with respect to the magnitude of the price change, and of course, a significant emphasis of our study is the possibility of mismatched signs between the true $GTE$ and the estimated $GTE$, which they do not study. In the appendix, we also show that these results generalize to a setting where a platform with heterogeneous goods for sale sets a fixed fee as a percentage of overall transaction volume, as is common in online marketplaces. Our results have important implications for price optimization based on experiments: following the suggestion of a canonical experiment may lead one to move in the wrong direction and decrease profits relative to the status-quo!
Taking advantage of our structural characterization of bias, in Section 6, we construct a novel debiasing method for canonical A/B experiments. In particular, we show that if the experimenter equally allocates units between treatment and control, it is possible to fully debias the canonical estimator by taking advantage of symmetry in the structure of bias. Notably, this allows firms to run the most simple version of an A/B pricing experiment, while still emerging with an unbiased estimator (under a set of assumptions that me make clear below). This is a contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require more complicated experimental designs and estimators (Holtz et al., 2020, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
To obtain additional insight, in Section 7, we apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022). We show how the biases of LR and CR experiments depend on model parameters, with a particular focus on regimes in which a “change of sign” may occur. For example, for an LR experiment, a platform is more likely to obtain the incorrect sign in an experiment if the demand is low relative to supply. We also investigate numerically how bias depends on both market balance and the baseline control price, and show how canonical A/B methods for price experimentation may result in significant profit underperformance.
In Section 8, we conclude the paper by calibrating our two-sided market model to empirical estimates from Farronato and Fradkin, (2022) for Airbnb markets in major U.S. cities. This exercise simplifies the Airbnb marketplace in many dimensions, but we believe it is useful for offering insights into the prevalence of estimators with the wrong sign in real-world markets. In our investigation, we find that such estimators are plausible in the Airbnb context. Notably, all markets in the study yielded estimators with the wrong sign when we simulated LR experiments, which would be a typical approach for testing pricing changes in practice. Although this analysis is based on a simplified market model, the results suggest that changes of sign are prevalent enough to be consequential and of concern for practitioners. We conclude in Section 9.
2 Related Work
The phenomenon we study is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts. More generally, this interference effect is a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Imbens and Rubin, 2015). Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014, Fradkin, 2019, Liu et al., 2020), showing its importance in practice. Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020, Candogan et al., 2024), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), rollout experiments Han et al., (2023), and message passing (Shirani and Bayati, 2024).
A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), Roemheld and Rao, (2024), Deng and Le, (2023), Simchi-Levi and Wang, (2023). Some are less related to our work; for example, Simchi-Levi and Wang, (2023) focuses on designing price experiments to mitigate the tail risk of revenue loss from experimentation. More related to our paper, both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects). Roemheld and Rao, (2024) uses an empirical study to suggest that false positives may be inflated due to interference in pricing experiments. More similar to our findings, Deng and Le, (2023) also notes the possibility of the sign of the treatment effect estimator being incorrect; in their setting this is due to an effect they call lead-day bias in markets with expiring goods. In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign. Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper—those where a continuous variable (such as price) is incrementally varied—as marginal policy changes, and develops a framework to study them.
The closest of these works to ours is Munro et al., (2021), which studies general estimands in an economic price equilibrium setting, and decomposes the treatment effect into direct and indirect effects; the former is obtained via “canonical” estimators, while the latter corresponds to our bias, and the authors show how it can be corrected using local price experiments. In that work, the direct and indirect effects are obtained as marginal changes with respect to the treatment assignment probability (inspired by work in Hu et al., (2022)), rather than the price as in our present paper. Further, our emphasis in applying our results is on structural insight regarding when changes of sign may be obtained for objectives such as profits or revenue, while their paper is focused on design and estimation for general estimands in market equilibrium. (An earlier version of Munro et al., (2021) considers treatment changes in a continuous variable, with an emphasis on relating the resulting differential treatment effect to direct and indirect effects.)
3 General Model Formulation
This section presents our general model, which we subsequently adapt to specific settings.
3.1 Market Model and Estimand
We consider a market in which a firm or platform sets a single variable of interest, $x\in\mathbb{R}_{+}$, evaluated by a performance metric $T(x)\in\mathbb{R},\forall x\in\mathbb{R}_{+}$. The performance metric can describe bookings, profits, or revenues of the platform. We make the following assumption that is kept throughout the paper. {assumption} The function $T(x)$ is differentiable for all $x\in\mathbb{R}_{+}$.
The firm is interested in optimizing the performance metric by adjusting the decision variable $x$. In particular, the firm is interested in the change in the target metric, when the decision variable takes on a new “treatment” value, denoted by $x_{1}$, moving away from the status-quo “control” value, denoted by $x_{0}$. This difference, $T(x_{1})-T(x_{0})$, corresponds to the global treatment effect ($GTE$). We define a local version of the $GTE$ as the derivative of the target metric, representing the case in which the treatment value of the decision variable is “close” to the control value:
$GTE(x_{0})$ $:=T^{\prime}(x_{0})\ .$ (1)
$GTE(x_{0})$ is our estimand.
3.2 Experiment: Design, Model, and Estimator
Our work focuses on settings where the function $T$ may be unknown to the firm or platform. Therefore, an experiment is conducted to determine the direction in which the decision variable should be adjusted in order to improve the performance metric (a form of policy improvement). We assume that the platform conducts an experiment with a single treatment condition where a fraction $q$ of the units are randomized to treatment and a fraction $1-q$ to control. We carry out our analysis in a formal “mean field” or continuum regime, where a continuum mass of units are allocated to each group; this allows us to focus on the magnitude and sign of bias (see also Johari et al., (2022), Wager and Xu, (2021) for related works using a similar approach).
It is worth noting that in the settings we are considering, randomization can occur on either side of the market, i.e., customers or products (or listings). In a customer-side randomization experiment (CR), a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition. On the other hand, in a listing-side randomization experiment (LR), a customer is exposed to a mix of treatment and control products (or listings). As before, we denote by $x_{0}$ the baseline (control) value of the decision variable and by $x_{1}$ the treatment value.
We denote the in-experiment performance metrics for control and treatment goods by the functions $T_{0}(x_{0},x_{1},q)$ and $T_{1}(x_{0},x_{1},q)$ respectively. The function $T_{0}(x_{0},x_{1},q)$ (resp., $T_{1}(x_{0},x_{1},q)$) corresponds to the performance metric in the control (resp., treatment) group when the control and treatment values are given by $x_{0}$ and $x_{1}$, respectively, and a fraction $q$ of units are randomized into treatment. The nature of these functions will become more clear in the context of the applications described below. Notably, this notation abstracts away from the specific experimental design (whether CR or LR). We make the following assumption that is kept throughout the paper: {assumption} The functions $T_{0}(x_{0},x_{1},q)$, and $T_{1}(x_{0},x_{1},q)$ are differentiable for all $x_{0},x_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
The canonical experimental designs at each side of the market have an associated natural “canonical” local estimator. To define this estimator, consider a sequence of experiments for which the treatment value of the decision variable approaches the control value. Then, we compute the relative difference between the performance metrics appropriately normalized by the treatment and control fractions, respectively, which corresponds to the Horvitz-Thompson estimator, or equivalently in our formal mean-field setting, the difference-in-means estimator. More formally, we consider a sequence of experiments indexed by $n\in\mathbb{Z}_{+}$, in which the treatment value approaches the control value: $\lim_{n\to\infty}x_{1}^{n}=x_{0}$. We define a local version of the canonical estimator as:
$\widehat{GTE}(x_{0},q)$ $:=\lim_{n\to\infty}\frac{\frac{T_{1}(x_{0},x_{1}^{n},q)}{q}-\frac{T_{0}(x_{0},x_{1}^{n},q)}{(1-q)}}{x_{1}^{n}-x_{0}}\ .$ (2)
We make the following assumption that is kept throughout the paper:
{assumption}
$T_{0}(x,x,q)/(1-q)=T_{1}(x,x,q)/q=T(x)$, for all $x\in\mathbb{R}_{+}$ and $10<q<$ 1.
The assumption states that if the values of control and treatment decision variables are the same, then both the scaled control and treatment performance metrics correspond to the market-wide metric. The assumption will play a key role in our analysis. To further understand it, consider $T_{1}$ with an extended domain $0<q\leq 1$ (the argument for $T_{0}$ is symmetric). Now, suppose $T_{1}$ is linear in its third argument over the relevant domain, that is, for any $0<q\leq 1$ and $10<\alpha<$ 1, $T_{1}(x,x,\alpha q)=\alpha T_{1}(x,x,q)$. If we take $q=1$, then $T_{1}(x,x,\alpha)/\alpha=T_{1}(x,x,1)=T(x)$, where the second equation follows directly by definition. Hence, the assumption is implied by constant (i.e., linear) returns of the performance metric in the fraction randomized. Note that in practice, experimental estimates typically implicitly rely on the assumption of constant returns in order to extrapolate estimates for control and treatment groups to the entire market. For the marketplace models and metrics we study later, this assumption will be satisfied.
Now, we obtain:
$\widehat{GTE}(x_{0},q)$ $=\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{1}(x_{0},x_{0},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$
$=\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{0}(x_{0},x_{0},q)/(1-q)-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$
$=T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\ .$ (3)
The first equation is direct from the definition of $\widehat{GTE}(x_{0},q)$, the second follows by Assumption 3.2, and the third by Assumption 3.2. The subscript $x$ (resp., $y$) denotes the partial derivative with respect to the first (resp., second) argument of the function.
Finally, we define the estimation bias of this local estimator as: [^3] Note that in statistics, bias is often defined as the estimator minus the estimand. However, we make the preceding choice because (as we will see) the resulting sign of $Bias$ is typically nonnegative, making it easier to obtain practical interpretations of our results.
$Bias(x_{0},q)$ $:=GTE(x_{0})-\widehat{GTE}(x_{0},q)$ (4)
$=T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)\ .$ (5)
Studying bias in this local regime proves fruitful, as it leads to the following simple characterization.
Proposition 3.1
The bias is given by the following expression:
$Bias(x_{0},q)=T_{0y}(x_{0},x_{0},q)/(1-q)+T_{1x}(x_{0},x_{0},q)/q\ .$
Proof 3.2
Proof. By Assumption 3.2, we have that:
$T(x_{0})=T_{1}(x_{0},x_{0},q)/q\ .$
By Assumption 3.2 and application of the chain rule we get:
$T^{\prime}(x_{0})=T_{1x}(x_{0},x_{0},q)/q+T_{1y}(x_{0},x_{0},q)/q\ .$
Replacing in (5), we obtain the desired result:
$Bias(x_{0},q)=T_{0y}(x_{0},x_{0},q)/(1-q)+T_{1x}(x_{0},x_{0},q)/q\ .$
$\Box$
The result provides an intuitive characterization of bias. Specifically, $T_{0y}(x_{0},x_{0},q)$ and $T_{1x}(x_{0},x_{0},q)$ correspond to the relative changes in performance for the control and treatment groups, respectively, when the decision variable of the other group changes marginally. The cross-partial derivatives represent the effect on the performance metric across experimental groups, making them an intuitive indicator of interference. As expected, as this interference effect increases, bias correspondingly increases.
4 Revenue Management Setting
In this section we examine a traditional revenue management setting with monopoly pricing. A firm or platform offers products that, at least based on observable characteristics, are deemed homogeneous and are thus priced uniformly. The firm faces an (unknown to them) demand function given by $D(p)$ as a function of the price $p$. The firm has a constant marginal production cost, $c\geq 0$. We make the following assumption that is kept throughout the paper and will imply Assumption 3.1 for the performance metrics we study.
{assumption}
The function $D(p)$ is differentiable and $D^{\prime}(p)<0$, for all $p\in\mathbb{R}_{+}$. At this level of abstraction, the demand function can describe a scenario with either an infinite supply or a limited supply, where a rationing mechanism determines the consumption. In the limit of a large market, even with supply constraints, we expect the demand function to be differentiable.
To provide a concrete example of a limiting regime with infinite supply, the demand function can be understood as the fraction among a large population of customers buying the product at a given price. That is, suppose $x_{i}(p)\in\{0,1\},i=1,2,\ldots$ are i.i.d. Bernoulli random variables with mean $0\leq D(p)\leq 1$; we interpret $x_{i}(p)$ as the purchasing decision of customer $i$ at price $p$. Then trivially by the law of large numbers $D(p)=\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}x_{i}(p)$.
4.1 Estimands
We will consider two performance metrics: (1) demand itself, which serves as a measure of growth, and (2) the firm’s profits (or revenue, in case costs are zero). The firm’s decision variable is price, and it is seeking to gain an understanding of how changes in price impact demand (i.e., price elasticity of demand), and subsequently affect profit. To achieve this goal, the firm is interested in estimating the difference in demand or profits generated when products are priced at a new “treatment” price relative to the demand or profits generated when products are priced at the status-quo “control” price. These differences correspond to the global treatment effects of interest that we derive below. To do so, we first define the “global” profit function as:
$\pi(p):=(p-c)D(p)\ .$
Following equation (3.1), we present our estimands. These are the local versions of the global treatment effects associated to demand and profits:
$GTE_{D}(p)$ $:=D^{\prime}(p)\ .$ (6)
$GTE_{\pi}(p)$ $:=\pi^{\prime}(p)$
$=D(p)+(p-c)D^{\prime}(p)\ .$ (7)
Note that $GTE_{\pi}(p)$ is the marginal profit of the firm at price $p$.
4.2 Experiment: Design, Model, and Estimator
Given that the demand function is unknown, we assume the firm uses price experimentation to estimate the $GTE$s of interest. As mentioned above, canonical A/B-tests can randomize on either side of the market, and our model does not preclude either possibility. That said, we do note that when conducting price experiments in practice, companies prefer to refrain from randomizing prices on the customer side to adhere to the basic principle that each customer should be offered the same price for the same product (Phillips, 2021).
We assume a single treatment condition in which a fraction $q$ of the subjects are randomized to treatment and a fraction $1-q$ to control. We denote by $p_{0}$ the baseline (control) price and by $p_{1}$ the treatment price. We assume that the experimental demands of control and treated subjects are given by the functions $Q_{0}(p_{0},p_{1},q)$ and $Q_{1}(p_{0},p_{1},q)$, respectively. We further define the functions:
$D_{0}(p_{0},p_{1},q)=\frac{Q_{0}(p_{0},p_{1},q)}{1-q};\quad D_{1}(p_{0},p_{1},q)=\frac{Q_{1}(p_{0},p_{1},q)}{q},$
which represent the demand for control and treated units scaled by the control and treatment fractions, respectively. Henceforth, we refer to $D_{0}$ and $D_{1}$ as the control and treatment demand functions, respectively. Again, at this level of abstraction these demand functions could, for example, incorporate supply constraints and rationing mechanisms.
Recall that in a CR experiment treatment and control customers will observe all products priced at the treatment and control prices respectively, with the proportion of customers assigned to the treatment group being $q$. In contrast, in a LR experiment, a customer will face an array of products, a fraction $q$ of which are priced at $p_{1}$ and the rest at $p_{0}$.
We make the following assumption that implies Assumption 3.2 and that is kept throughout the paper. {assumption} The functions $D_{0}(p_{0},p_{1},q)$, and $D_{1}(p_{0},p_{1},q)$ are differentiable for all $p_{0},p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1. In addition, $p_{0},p_{1}\geq c$. Note that the assumption precludes the case of perfect substitution between control and treatment products. Under perfect substitution, a negligible undercut of the treatment relative to the control price would result in a discrete increase of the treatment demand; as a result, the demand function would not be continuous. Even though, as initially stated, the firm is selling a single type of product, we can rationalize the smoothness assumption in at least one of two ways, commonly assumed in pricing settings: (1) products can be identical in terms of observable characteristics but there may be latent horizontal differentiation; (2) there could be search frictions that do not allow customers to always buy the lowest-priced product.
We introduce the following additional assumption on the experimental demand functions which implies Assumption 3.2. {assumption} The following statements all hold:
  1. 1.
    $D_{0}(p_{0},p_{1},q)$ is non-decreasing in $p_{1}$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
  2. 2.
    $D_{1}(p_{0},p_{1},q)$ is non-decreasing in $p_{0}$, for all $p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
  3. 3.
    $D_{0}(p,p,q)=D_{1}(p,p,q)=D(p)$, for all $p\in\mathbb{R}_{+}$ and $10<q<$ 1.
The first two statements in Assumption 4.2 are natural as they imply that control and treatment products are (imperfect) substitutes. The third assumption states that if the control and treatment prices are the same, then both the scaled control and treatment demand functions are equivalent to the market demand function; this assumption is implied by constant returns to scale. This assumption is satisfied by demand systems for which the control and treatment demand quantities are proportional to their shares in the marketplace. Indeed, this assumption will be satisfied for the model we later study in Section 7, as well for a number of other common demand systems.
We finish this subsection by defining the experimental estimators. To do so, we define the experimental control and treatment profit functions, respectively:
$\pi_{0}(p_{0},p_{1},q)$ $:=(p_{0}-c)D_{0}(p_{0},p_{1},q)\ ,$
$\pi_{1}(p_{0},p_{1},q)$ $:=(p_{1}-c)D_{1}(p_{0},p_{1},q)\ .$
Following Definition (2) we define canonical estimators for demand and profit associated with the canonical experimental designs:
$\widehat{GTE}_{D}(p_{0},q)$ $:=\lim_{n\to\infty}\frac{D_{1}(p_{0},p_{1}^{n},q)-D_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}\ ,$
$\widehat{GTE}_{\pi}(p_{0},q)$ $:=\lim_{n\to\infty}\frac{\pi_{1}(p_{0},p_{1}^{n},q)-\pi_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}\ .$
We note that analogs of these canonical estimators are frequently used in practice in estimating global treatment effects from experimental data; in particular, in our continuum model, these estimators naturally correspond to difference-in-means or Horvitz-Thompson type estimators, scaled by the price difference to estimate the target estimand (the derivative of demand or profit at the control price).
Now, by equation (3), we obtain:
$\widehat{GTE}_{D}(p_{0},q)$ $=D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\ ,$ (8)
$\widehat{GTE}_{\pi}(p_{0},q)$ $=\pi_{1y}(p_{0},p_{0},q)-\pi_{0y}(p_{0},p_{0},q)$
$=D(p_{0})+(p_{0}-c)\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\ ,$ (9)
where we used Assumption 4.2 and the last equation follows by Assumption 4.2.3.
5 Biases: Analysis and Change of Sign
In this section we analyze the biases associated with the canonical estimators introduced above. Following Definition (4), we define the estimation biases for demand and profit as:
$Bias_{D}(p_{0},q)$ $:=GTE_{D}(p_{0})-\widehat{GTE}_{D}(p_{0},q)\ ,$
$Bias_{\pi}(p_{0},q)$ $:=GTE_{\pi}(p_{0})-\widehat{GTE}_{\pi}(p_{0},q)\ .$
First, note that by Proposition 3.1 we obtain the following bias characterizations:
$Bias_{D}(p_{0},q)$ $=D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\ ,$ (10)
$Bias_{\pi}(p_{0},q)$ $=\pi_{0y}(p_{0},p_{0},q)+\pi_{1x}(p_{0},p_{0},q)$
$=(p_{0}-c)\left(D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\right)$ (11)
$=(p_{0}-c)Bias_{D}(p_{0},q)\ .$ (12)
As discussed in Proposition 3.1, bias increases as interference effects between treatment and control groups increase. More specifically, the bias is influenced by the degree to which the control (or treatment) group cannibalizes demand from its counterpart in response to a marginal price change in the latter group.
Now, note that under Assumptions 4.2 and 4.2, both $Bias_{D}(p_{0},q)$ and $Bias_{\pi}(p_{0},q)$ are non-negative, meaning the canonical estimators always (weakly) underestimate the $GTE$:
Proposition 5.1
We have that $\widehat{GTE}_{D}(p_{0},q)\leq GTE_{D}(p_{0})$ and $\widehat{GTE}_{\pi}(p_{0},q)\leq GTE_{\pi}(p_{0})$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
This result follows similar cannibalization arguments developed in previous work (Johari et al., 2022). To provide intuition, consider an LR experiment in which the treatment price is higher than control the price. In this case, experimental treatment demand underestimates global treatment demand because of substitution in the experiment from treatment units to control units. Similarly, experimental control demand overestimates global control demand. Instead, let’s consider a CR experiment where there are supply constraints, so customers compete for limited products. In such cases, the control customers may “cannibalize” supply from the treatment group, because of their lower control price, causing an underestimation of the global treatment effect again. Note that in a CR experiment without supply constraints, we anticipate no bias because there is no competition between the treatment and control groups.
It is important to observe that because $D^{\prime}(p)<0$ by Assumption 4, $GTE_{D}(p_{0})<0$, for all $p$. Moreover, since $\widehat{GTE}_{D}(p_{0},q)$ is an underestimation, both the estimand and the canonical estimator will have the same (nonpositive) sign.
However, there is a more intriguing result for profits. Suppose that the treatment price $p_{1}$ is greater than the control price $p_{0}$. In this scenario, even if the demand $D(p_{1})$ at $p_{1}$ is lower than the demand $D(p_{0})$ at $p_{0}$, the increase in price may offset the decrease in demand, leading to a positive true global treatment effect in profits. But, because the canonical estimator underestimates the treatment effect, the profit GTE estimator could be negative. Next, we will provide conditions under which the canonical estimator not only underestimates the global treatment effect, but actually obtains the wrong sign.
By definition, $GTE_{\pi}(p_{0})\geq 0$ if and only if $\pi^{\prime}(p_{0})\geq 0$, which after some simple manipulation yields the following standard “markup” inequality:
[…]
where $e_{p}(p_{0})=D^{\prime}(p_{0})p_{0}/D(p_{0})$ is the price elasticity of demand. If $\pi(p)$ is strictly quasi-concave, the condition is equivalent to the price $p_{0}$ being below the optimal monopoly price. For the case $c=0$, condition (a) naturally states that marginal revenue is increasing at $p_{0}$, or equivalently that the demand is relatively inelastic with respect to price.
Now, consider the canonical estimator. Note that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ is equivalent to $Bias_{\pi}(p_{0},q)\geq{GTE}_{\pi}(p_{0})$. That is, provided that $GTE_{\pi}(p_{0})\geq 0$ holds, we have a change of sign when the bias is so severe that it becomes larger than the treatment effect. Furthermore, after some manipulation of equation (9), we obtain that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ if and only if:
[…]
Note that condition (b) is a “modified markup” inequality. We have the following result.
Proposition 5.2
  1. 1.
    Condition (a) holds if and only if $GTE_{\pi}(p_{0})\geq 0$.
  2. 2.
    Condition (b) holds if and only if $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$.
  3. 3.
    $A(p_{0})\leq B(p_{0},q)$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
Proof 5.3
Proof. Points (1) and (2) in the proposition are direct from the arguments above. To prove point (3), note that $D_{1y}(p_{0},p_{0},q)\leq D^{\prime}(p_{0})$ (by applying Assumptions 4.2.2 and 4.2.3 together with the chain rule) and $D_{0y}(p_{0},p_{0},q)\geq 0$ (by Assumption 4.2.1). Hence,
$\frac{p_{0}-c}{p_{0}}\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\leq\frac{p_{0}-c}{p_{0}}D^{\prime}(p_{0})\ ,$
and thus the result follows because $D^{\prime}(p_{0})<0$ by Assumption 4. $\Box$
The proposition states a sign change is obtained, $\widehat{GTE}_{\pi}(p_{0},q)\leq 0\leq GTE_{\pi}(p_{0})$, if and only if conditions (a) and (b) hold. This implies lower and upper bounds on the inverse of the price elasticity of demand. Note that we use weak inequalities in conditions (a) and (b) for a change of sign; strictly speaking of course, a change of sign can only occur if at least one of the two inequalities is strict. [^4] If exactly one of these inequalities is strict (and both hold), the canonical estimator will still instruct firms to move in a non-optimal direction; either by telling them to lower prices when they were already at the optimal price, or by suggesting they keep their price constant when they were underpricing. We suppress this edge case in the sequel in favor of simplicity of presentation. [^5] In our two-sided market model, introduced in section 7, the scenario $\widehat{GTE}_{\pi}(p_{0},q)=GTE_{\pi}(p_{0})=0$ is impossible, meaning this edge case does not exist.
To provide intuition, note that by using equation (8) we can define the “experimental” price elasticity of demand as:
$\widehat{e_{p}}(p_{0},q)=\frac{p_{0}(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q))}{D(p_{0})}.$ (15)
Hence, we can re-write condition (b) as:
$-\frac{1}{\widehat{e_{p}}(p_{0},q)}\leq\frac{(p_{0}-c)}{p_{0}}\ ,$
which can be understood as the experimental analogue of a markup inequality. It states that the experimental elasticity suggests the current price is above the optimal monopoly price (under an appropriate concavity assumption). Hence, as expected, a change of sign occurs when the canonical estimator indicates that the platform is overpricing relative to the optimal monopoly price (and therefore prefers to decrease prices), when in reality it is underpricing (and therefore should raise prices). This result has important consequences for experiment-driven price optimization: the prescription implied an the experiment (reduce price) may be the opposite of the one implied by the $GTE$ (increase price). Therefore, if a firm moves in the direction suggested by the experiment, profits have the potential to decrease relative to the status-quo.
To provide further intuition, let us assume that condition (a) holds and that profits are quasiconcave, so that the firm is underpricing relative to the optimal monopoly price. Hence, we have a change of sign if and only if condition (b) holds. If $|D_{1y}(p_{0},p_{0},q)|+D_{0y}(p_{0},p_{0},q)$ is large compared to $|D^{\prime}(p_{0})|$, condition (b) will hold. Note that $|D_{1y}(p_{0},p_{0},q)|$ represents the magnitude of the decrease in treatment demand given a small increase in treatment price, which captures two effects: (i) demand lost to the outside option; and (ii) cannibalization to the control group. Note that (i) is the global demand effect we aim to estimate. On the other hand, the term $D_{0y}(p_{0},p_{0},q)$ represents the increase in control demand given a small increase in treatment price, which captures cannibalization from the treatment to the control group (equivalent to effect (ii)). A change of sign happens as the two cannibalization effects just mentioned become larger.
We conclude with two remarks regarding our estimators, our findings on changes of sign, and guidance for practitioners. First, note that the estimators $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that we defined are estimates of derivatives of the demand and profit functions, respectively, as is natural given our estimands. In practice, a platform may use a difference in estimated treatment and control demand, or estimated treatment and control profit, without normalizing by the difference in prices. This can have some bearing in interpreting our results, depending on the sign of the difference between the treatment and control price.
Concretely, suppose that treatment involves raising the price relative to control. Then our results show that a change of sign can only occur if the unnormalized global treatment effect on profits is positive, but the treatment effect is estimated to be negative in the experiment. On the other hand, suppose the treatment involves lowering the price relative to control. Then our results show that a change of sign can only occur if the unnormalized global treatment effect on profits is negative, but the treatment effect is estimated to be positive in the experiment. In either case, the platform risks either failing to move prices in a profitable direction, or worse, moving them in an unprofitable direction.
Second, we note that in our case, we have studied a model where the platform changes the price directly. In practice, another relevant platform parameter that affects profits can be a platform-wide fee. Specifically, in Appendix 12 , we present a model where a platform offers a multitude of heterogeneous goods (potentially at different prices), and earns profit by charging a percentage fee on each transaction. This is a common monetization model for online marketplaces. Viewing the fee as the variable of interest for the platform’s profit optimization, we obtain analogous results to those in this section, including the bias characterization, the possibility of a change of sign, and interpretation of the conditions for a change of sign in terms of appropriate derivatives . This extension suggests our insights in this paper also have implications for practitioners in platform settings with many heterogeneously differentiated goods, where fees are the key variable for monetization.
6 Debiasing Canonical Estimators
We take advantage of our structural characterization of bias to introduce a novel approach to debiasing $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$. The key to this approach, and a necessary step for it to be applied, is for the experimenter to equally allocate units between treatment and control, i.e., setting $q=1/2$.
Recall that equations (10) and (11) quantify bias in terms of the derivatives of demand. We will now show how we can use observables to estimate this bias. First, we show that $D_{0y}(p_{0},p_{0},q)$ can be computed (approximately) using observed data. To see this, note that $D_{0y}(p_{0},p_{0},q)$ is defined by:
$D_{0y}(p_{0},p_{0},q)=\lim_{p_{1}\to p_{0}^{+}}\frac{D_{0}(p_{0},p_{1},q)-D_{0}(p_{0},p_{0},q)}{p_{1}-p_{0}}.$
By Assumption 6 we have that $D_{0}(p_{0},p_{0},q)=D(p_{0})$. To make progress, we make the assumption that since the firm was pricing at $p_{0}$ before the experiment began, the pre-experiment observation of demand is a reasonable proxy for the control demand $D(p_{0})$. The quantity $D_{0}(p_{0},p_{1},q)$ is observed as well: it represents the control demand during the experiment. This means that in the limit as $p_{1}$ approaches $p_{0}$, observables can be used to approximately compute the derivative $D_{0y}(p_{0},p_{0},q)$.
We now note that by symmetry, we have $D_{0}(p_{0},p_{1},1/2)=D_{1}(p_{1},p_{0},1/2)$. This holds because both terms represent demand for the half of the goods that are priced at $p_{0}$, when the other half of goods are priced at $p_{1}$. Taking the derivative of both sides of this equality with respect to $p_{1}$, we obtain $D_{0y}(p_{0},p_{1},1/2)=D_{1x}(p_{1},p_{0},1/2)$, so that when $p_{1}=p_{0}$, there holds $D_{0y}(p_{0},p_{0},1/2)=D_{1x}(p_{0},p_{0},1/2)$. We have thus shown the following result.
Proposition 6.1
Suppose that $q=1/2$. Then there holds:
$Bias_{D}(p_{0},1/2)$ $=2D_{0y}(p_{0},p_{0},1/2);$ (16)
$Bias_{\pi}(p_{0},1/2)$ $=2(p_{0}-c)D_{0y}(p_{0},p_{0},1/2).$ (17)
In particular, there holds:
$GTE_{D}$ $=\widehat{GTE}_{D}(p_{0},1/2)+2D_{0y}(p_{0},p_{0},1/2);$ (18)
$GTE_{\pi}$ $=\widehat{GTE}_{\pi}(p_{0},1/2)+2(p_{0}-c)D_{0y}(p_{0},p_{0},1/2).$ (19)
Thus if the firm is able to estimate $D_{0y}(p_{0},p_{0},q)$ using observables (as described above), then the right hand sides of (18) and (19) are debiased estimators of $GTE_{D}$ and $GTE_{\pi}$ respectively.
We note three limitations of this method. First, of course, the symmetry property being exploited only applies when the market is equally allocated to treatment and control. Second, the method is not sample efficient: only half the experimental samples, those from the control group, are used. Finally, our view that the platform can accurately estimate $D(p_{0})$ depends on the market remaining reasonably stationary during the experiment; if the market is nonstationary, this adjusted estimator may not capture the true treatment effect during the experiment.
7 A Two-Sided Market Model
In this section, we illustrate the application of our results from the general revenue management setting (introduced in Section 4) to the two-sided market model developed in Johari et al., (2022). That paper develops a stochastic dynamic market model with supply constraints in which customers book listings for a finite period of time. The paper also introduces a mean-field limit that greatly simplifies the analysis. We focus on the mean-field model here and describe its main elements.
7.1 Preliminaries
We begin by recapping the mean-field model formulation in Johari et al., (2022), adapted to our setting with price experiments. There is a two-sided platform evolving in continuous time with listings on the supply side and customers on the demand side. There is a continuum of listings of mass $\rho$ that are homogenous with respect to their observable characteristics. Having a limited supply will play a key role in the analysis. We assume all listings have the same price $p$. At each time $t$, listings can be available or booked. The state of the system at time $t$ is given by $s_{t}$, which denotes the mass of available listings available at time $t$.
[…]
$\frac{sv(p)}{\varepsilon+sv(p)}.$
Finally, a listing remains occupied for an expected time $\frac{1}{\tau}$ when it gets booked; in our analysis we normalize $\tau$ to be 1 without loss of generality. Note that in our model, the parameter $\lambda$ mediates market balance: when $\lambda\rightarrow 0$ the market becomes demand constrained, when $\lambda\rightarrow\infty$ the market becomes supply constrained.
The general model in Johari et al., (2022) allows for heterogeneity in both listings and consumers. To simplify the analysis and exposition, in the model we consider in this section, we make both consumers and listings homogeneous.
As in Johari et al., (2022) we focus on the steady-state of the system. Let $s^{*}(\lambda,p)$ be the steady-state mass of available listings for a given market balance parameter $\lambda$ and price $p$. It is shown in Johari et al., (2022) that $s^{*}(\lambda,p)$ is the unique solution to the following conservation of flow equation:
$\rho-s^{*}(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)v(p)}{\varepsilon+s^{*}(\lambda,p)v(p)}:=D(\lambda,p)\ .$ (20)
The left-hand side of the equation corresponds to the rate at which listings become available. The right-hand side corresponds to the booking (or demand) rate, defined as $D(\lambda,p)$. In the steady-state both rates are equal.
{assumption}
We assume the following for the utility function $v(p)$:
  1. 1.
    $v(p)$ is strictly positive for $p\in[c,+\infty);$
  2. 2.
    $v(p)$ is strictly decreasing for $p\in[c,+\infty);$
  3. 3.
    $v(p)$ is differentiable for $p\in[c,+\infty)$, and $v^{\prime}(p)$ is continuous;
  4. 4.
    $-pv^{\prime}(p)/v(p)$ is strictly increasing in $p$ for $p\in[c,+\infty).$
[…]
7.2 Steady-State and Estimand
To gain some initial intuition, we start this section by providing some analysis on the steady-state and the associated demand. Then, we define our estimands.
First, we note that listing availability $s^{*}(\lambda,p)$ is monotonically increasing with price $p$ and decreasing with the market balance parameter $\lambda$, which is natural. These monotonicity results can be derived by differentiating $s^{*}(\lambda,p)$ in equation (20). [^9] Indeed, we obtain $s^{*}_{p}(\lambda,p)=-\frac{s^{*}(\lambda,p)v^{\prime}(p)\varepsilon}{\frac{1}{\lambda}(\varepsilon+s^{*}(\lambda,p)v(p))^{2}+v(p)\varepsilon}>0$ and $s^{*}_{\lambda}(\lambda,p)=-\frac{v(p)s(\lambda,p)(v(p)s(\lambda,p)+\varepsilon)}{\varepsilon v(p)(2s(\lambda,p)+\lambda)+v(p)^{2}s(\lambda,p)^{2}+\varepsilon^{2}}<0$. Figure 6 in the Appendix illustrates the steady-state listing availability as a function of price ($p$) and market balance ($\lambda$).
Naturally, $D(\lambda,p)$ is monotonically decreasing in price ($p$) and increasing in market balance ($\lambda$). Again, we can prove these monotonicity results by differentiating. [^10] This can be seen by rewriting the demand as the rate at which the listings become available. Using the market balance equation, we have $D(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)v(p)}{\varepsilon+s^{*}(\lambda,p)v(p)}=\rho-s^{*}(\lambda,p)$. Hence, $D_{p}(\lambda,p)=-s^{*}_{p}(\lambda,p)$ and $D_{\lambda}(\lambda,p)=-s^{*}_{\lambda}(\lambda,p)$. Since $s^{*}_{p}(\lambda,p)>0$ and $s^{*}_{\lambda}(\lambda,p)<0$ we obtain that $D_{p}(\lambda,p)<0$ and $D_{\lambda}(\lambda,p)>0$. Figure 7 in the Appendix shows the dependence of the demand function on the market balance parameter and price.
Now, we define our estimands. Following the definitions in equations (6) and (7):
$GTE_{D}(\lambda,p)$ $=D_{p}(\lambda,p)$
$GTE_{\pi}(\lambda,p)$ $=D(\lambda,p)+(p-c)D_{p}(\lambda,p)$
$=\rho-s^{*}(\lambda,p)-(p-c)s^{*}_{p}(\lambda,p),$ (21)
where in the last equation we used equation (20).
7.3 Experimental Designs
In this section we analyze two canonical experimental designs in platforms: randomization across listings (LR experimentation) and randomization across customers (CR experimentation).
7.3.1 LR Experimentation
We assume that randomization occurs on the listing side, where a fraction $q$ of the listings are assigned to treatment. Hence, a customer is exposed to a mix of treatment and control listings. In this experimental design, the balance equations are given by:
$(1-q)\rho-s_{0}^{*}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s_{0}^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s_{0}^{*}(\lambda,p_{0},p_{1},q)v(p_{0})+s_{1}^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}$ (22)
$q\rho-s_{1}^{*}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s_{1}^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s_{0}^{*}(\lambda,p_{0},p_{1},q)v(p_{0})+s_{1}^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}\ ,$
where $s^{*}_{0}(\lambda,p_{0},p_{1},q)$ and $s^{*}_{1}(\lambda,p_{0},p_{1},q)$ refer to the steady-state available control and treated listings, respectively.
The control and treatment demand functions $D^{LR}_{0}(\lambda,p_{0},p_{1},q)$ and $D_{1}^{LR}(\lambda,p_{0},p_{1},q)$ are given by:
$D_{0}^{LR}(\lambda,p_{0},p_{1},q)$ $=\rho-s_{0}^{*}(\lambda,p_{0},p_{1},q)/(1-q);$ (23)
$D^{LR}_{1}(\lambda,p_{0},p_{1},q)$ $=\rho-s_{1}^{*}(\lambda,p_{0},p_{1},q)/q.$
We can show that $D^{LR}_{0}(\lambda,p_{0},p_{1},q)$ and $D^{LR}_{1}(\lambda,p_{0},p_{1},q)$ satisfy Assumption 4.2 (see Appendix 15.2).
Following the characterization of bias from equations (10) and (12), we obtain that the bias in this setting, $Bias_{\pi}^{LR}(\lambda,p_{0},q)$ , is given by:
$Bias^{LR}_{\pi}(\lambda,p_{0},q)$ $=(p_{0}-c)(D^{LR}_{1x}(\lambda,p_{0},p_{0},q)+D^{LR}_{0y}(\lambda,p_{0},p_{0},q))$
$=(p_{0}-c)\frac{-\lambda s^{*}(\lambda,p_{0})^{2}v(p_{0})v^{\prime}(p_{0})(\varepsilon+s^{*}(\lambda,p_{0})v(p_{0}))}{[\varepsilon+s^{*}(\lambda,p_{0})v(p_{0})+\lambda v(p_{0})][(s^{*}(\lambda,p_{0})v(p_{0})+\varepsilon)^{2}+\lambda v(p_{0})\varepsilon]}$ (24)
Note that the bias is independent of the treatment fraction $q$.
7.3.2 CR Experimentation
In this setting, we randomize over customers, with a fraction $q$ of customers being placed into treatment, and the remainder in control. In this case, a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition. The steady-state is given by:
$\rho-s^{*}(\lambda,p_{0},p_{1},q)=\left[q\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}+(1-q)\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}\right].$ (25)
Observe that the two terms on the right hand side above are rescaled versions of the control and treatment demand functions:
$D^{CR}_{0}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})},$ (26)
$D^{CR}_{1}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}.$
One can show that the demand functions $D^{CR}_{0}(\lambda,p_{0},p_{1},q)$ and $D^{CR}_{1}(\lambda,p_{0},p_{1},q)$ satisfy Assumption 4.2 (see Appendix 15.3).
Using the characterization in equation (12) we obtain the following expression for $Bias_{\pi}^{CR}$:
$Bias^{CR}_{\pi}(\lambda,p_{0},q)$ $=(p_{0}-c)(D^{CR}_{1x}(\lambda,p_{0},p_{0},q)+D^{CR}_{0y}(\lambda,p_{0},p_{0},q))$
$=-(p_{0}-c)\frac{\lambda^{2}\varepsilon^{2}v(p_{0})v^{\prime}(p_{0})s^{*}(\lambda,p_{0})}{(v(p_{0})s^{*}(\lambda,p_{0})+\varepsilon)^{2}\left[(s^{*}(\lambda,p_{0})v(p_{0})+\varepsilon)^{2}+\lambda v(p_{0})\varepsilon\right]}$ (27)
Note that the bias in the CR-setting is independent of the treatment fraction $q$.
7.4 Properties of Biases with Market Balance
Johari et al., (2022) studies the bias of CR and LR experiments as the market becomes heavily demand constrained or supply constrained, respectively. Following the analysis of that paper, we obtain analogous insights, with proofs similar to Johari et al. (2022) but also leveraging our characterization of biases through cross-partial derivatives. In particular, as the market becomes more demand constrained ($\lambda\rightarrow 0$), competition between customers declines, resulting in less interference in a CR experiment; in the limit, the bias in a CR experiment approaches zero. On the other hand, in the same regime the competition between listings is high, resulting in more interference for an LR experiment, so in the limit the bias of an LR experiment remains positive. By contrast, as the market becomes more supply constrained ($\lambda\to\infty$), the opposite results hold: in this regime, competition between customers is stronger and competition between listings is weaker, so in the limit the bias of a CR experiment remains positive while the bias of a LR experiment approaches zero. We derive these results in Propositions 13.1 and 13.2 in Appendix 13.
Below we provide two figures to complement our limit results. In Figures 1 and 2 we present bias normalized by the arrival rate $\lambda$. We normalize by $\lambda$ instead of the GTE because the GTE for profits converges to 0 around the optimal price. As $\lambda$ increases, the market becomes more supply-constrained and bias for a LR experiment decreases as competition between listings dissipates. Figure 2 shows the normalized bias of demand and profits in a CR experiment and this dependence on $\lambda$ is naturally reversed. Both figures also show interesting behavior of biases with respect to the baseline price $p$, a novel aspect of our setting relative to Johari et al., (2022). As the price $p$ changes, the sensitivity of demand to price changes as well. In particular, in both CR and LR experiments, both $D_{0y}$ and $D_{1x}$ tend to be smaller at larger prices (i.e., demand becomes more inelastic as the price becomes large); this behavior drives the shape of the bias curve with respect to price, and subsequently profit. Note in particular that LR experiments exhibit interesting behavior when prices are large: despite the fact that we nominally expect bias to increase when the market is demand constrained, and higher prices reduce demand in the market, nevertheless the bias drops when prices are high. This is because of the fact that elasticities change with price, and illustrates the nuanced nature of interference in price experiments.
Figure 1: Illustration of $Bias^{LR}_{D}/\lambda$ and $Bias^{LR}_{\pi}/\lambda$ as a function of price ($p$) and market balance ($\lambda$). Instance is given by $v(p)=e^{V-p}$, $\epsilon=1$, $c=1$, $V=5$, and $\rho=1$.
Figure 2: Illustration of $Bias^{CR}_{D}/\lambda$ and $Bias^{LR}_{\pi}/\lambda$ as a function of price ($p$) and market balance ($\lambda$). Instance is given by $v(p)=e^{V-p}$, $\epsilon=1$, $c=1$, $V=5$, and $\rho=1$.
7.5 Change of Sign
We recall the two conditions for a change of sign to occur:
[…]
Interestingly, observe that the bias terms derived in Section 7.4 do not depend on the treatment fraction $q$ for either LR or CR experiments. In other words, the treatment fraction is not a lever that can be used to prevent—or even reduce—the possibility of a change of sign in the treatment effect estimator.
Using the structure of the model we can provide the following insightful characterizations for the change of sign regions.
Theorem 7.1
For all $\lambda\in(0,\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{LR}_{\lambda}\subseteq[c,\infty)$ such that the sign of an LR experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p\in[c,\infty)$ will be wrong if and only if $p\in P^{LR}_{\lambda}$.
Theorem 7.2
For all $\lambda\in(0,\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{CR}_{\lambda}\subseteq[c,\infty)$ such that the sign of a CR experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p\in[c,\infty)$ will be wrong if and only if $p\in P^{CR}_{\lambda}$.
Theorems 7.1 and 7.2 state that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$, such that the sign of the canonical estimator will be wrong, regardless of whether an LR or CR experiment is run. The proofs are provided in Appendix 15.7 and repeatedly exploit the market balance equations for CR and LR respectively.
The right hand side of both the intervals $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ is the optimal monopoly price $p^{*}(\lambda)$, i.e., where $GTE_{\pi}(\lambda,p^{*})=0$, for each $\lambda$. In particular, this reveals that the sets $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not comprised of unrealistic prices: if a firm is currently pricing optimally or very slightly below optimally, the sign of their canonical estimator will be wrong, regardless of their experimental setup.
It is not true, however, that the two experimental setups are equally fallible in all settings. Note that Condition (b) in 14 is different for LR and CR because the elasticities and demand sensitivities differ, meaning that the left endpoints of $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not the same. As expected based on the results of the previous section, they behave quite differently as a function of market balance, as we illustrate through the following numerical study.
7.5.1 Numerical Results
Figure (3) shows numerical results on the set of prices where a change of sign occurs.
Figure 3: Illustration of the change of sign region (purple) in an LR-experiment (left) and CR-experiment (right), respectively, with $v(p)=e^{V-p}$, $V=5$, $c=0.5$, $\rho=1$, and $\epsilon=1$. The vertical axis is the (log of) the arrival rate, or market balance, paramter $\lambda$; the horizontal axis is the (baseline) control price $p$, at which the experiment is conducted. The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red. The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in blue.
Consistent with Propositions 13.1 and 13.2, in Figure 3 we observe that the change of sign region becomes very small in the demand-constrained regime for a CR experiment and in the supply constrained regime for an LR experiment. At intermediate market balance regimes, we can more immediately observe that the change of sign region corresponds to a single continuous interval on prices, as Theorems 7.1 and 7.2 suggest. We observe that in the demand-constrained regime Condition (b) for a change of sign is more likely to be satisfied by the LR experiments, while in the supply-constrained regime Condition (b) is more likely to be satisfied by the CR experiments as expected. We also see that the region where a change of sign occurs is, overall, increasing with market balance for the CR experiments and decreasing for the LR experiment.
7.5.2 Gap from Optimality
One important implication of Figure 3 is the impact that the change of sign region has on the net profits of firms who use canonical price experimentation methods. Consider the practical importance of the change of sign region’s borders. The right border of the change of sign region (where purple meets red), represents the optimal price for that particular market balance parameter $\lambda$. This is the point where $GTE_{\pi}=0$. The left border of the change of sign region (where blue meets purple) represents the final price a firm would reach, were it to repeatedly update it’s price in response to canonical A/B experimentation. This is the point where $\widehat{GTE}_{\pi}=0$, or equivalently, where $Bias_{\pi}=GTE_{\pi}$. Note that this price corresponds to the optimum suggested by repeated canonical A/B experiments. This is because, to the right of that price point, the expriment suggests decreasing the price, and to the left, to increase it.
The difference between these two prices (the true optimal price and the experimentally optimized price) can correspond to significant losses in overall profit. In Figure 4, we see the percent of true optimal profit that a firm will achieve if it uses canonical LR (left) or CR (right) experimentation, for a range of market balance scenarios. What we find is that the profit achieved from optimization via experimentation can be far below the optimal profit, with the impact worsening in highly demand-constrained markets (for LR experiments) or demand-constrained markets (for CR experiments). Even when markets are reasonably balanced (i.e., $\lambda\approx 1$), both LR and CR experimentation will lead to a profit that is more than 4% lower than optimal. Overall, these findings suggest that firms who use canonical A/B methods for price experimentation might be operating well below their potentially optimal profits, even if they are in well balanced markets. Furthermore, these losses can increase significantly as markets become more imbalanced, getting to $20-40\%$ of optimal.
Figure 4: Illustration of the gap from optimal profits that canonical A/B experimentation creates.
8 An Empirical Case Study
Thus far, we have shown that A/B pricing experiments can get the sign wrong, and that associated profit losses can be significant in a two-sided market model with logit demand. To align our analysis even more closely with practical scenarios, we empirically calibrate our model from the previous section using published data from several Airbnb markets, and investigate the conditions under which the sign-change phenomenon may occur in price experiments within this context. Although our model simplifies the complexities of the Airbnb marketplace in many dimensions, we believe this exercise is valuable because it offers practical insights into how our results can align with real-world experiments. As will be shown in this section, our results suggest that a change of sign is indeed quite plausible in the empirical Airbnb context we study, demonstrating that these errors are not just a knife-edge issue—and thus potentially of high consequence for practitioners.
8.1 Model Calibration
In this section, we calibrate a specific version of our model from Section 7 to several Airbnb markets using the data from Farronato and Fradkin, (2022). Specifically, we consider the model from that section with $v(p)=e^{V-p}$. Note that this is a natural choice implied by a logit random utility model, where utility for the product is quasilinear in the price. For fixed $\varepsilon$, $V$, $\lambda$, and $p$, the resulting market balance equation is:
$\rho-s^{*}(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)e^{V-p}}{\varepsilon+s^{*}(\lambda,p)e^{V-p}}.$ (28)
Recall that then in this model, $D(p)=\rho-s^{*}(\lambda,p)$, and for a given supplier cost $c$, its profit is $\pi(p)=(p-c)D(p)$.
In this exercise, we consider that $GTE$ corresponds to suppliers’ profits and that the experiment is on actual prices. It is worth noting that in practice, Airbnb does not actually set the price of listings, but instead sets fees. However, we have elected to use our price testing framework (in place of the fee-model detailed in Appendix 12) because the estimates made in Farronato and Fradkin, (2022) are better suited to examining price experiments. . Airbnb actively provides pricing guidance to hosts aimed at optimizing their pricing strategies and regularly conducts A/B tests to evaluate these recommendations (Deng and Le, 2023). This setting aligns closely with the price testing framework employed in this section. Furthermore, since fee changes are often passed on to consumers through adjustments to prices, this is a reasonable approach for our analysis.
The goal of our calibration process is to arrive at imputed values of the parameters that define the model: (1) the total mass of listings $\rho$; (2) the outside option value $\varepsilon$; (3) the cost $c$; (4) the value $V$; (5) the arrival rate $\lambda$; and (6) the reference price $p$, from which the platform runs a price experiment.
We now describe how we arrive at these six parameters. We leverage structural estimates made by Farronato and Fradkin, (2022) for Airbnb markets in major U.S. cities. In particular, Farronato and Fradkin, (2022) segment the Airbnb market into four tiers: economy, midscale, upscale, and luxury. They study each of these tiers across 10 cities. We refer to each pair of tier and city as a segment, and use superscript $s$ to denote a segment. For each segment $s$, Farronato and Fradkin, (2022) reports (1) an estimate of the average cost (per night booked), which we denote $\hat{c}^{(s)}$; and (2) an estimate of own price elasticity of demand, which we denote $\hat{e}_{p}^{(s)}$. [^11] These values come directly from tables E8 and E14 in the Appendix of Farronato and Fradkin, (2022).
For our empirical study, we treat each segment as its own instantiation of the market model in (28). For example, if we are currently considering midscale listings in Detroit, then $v(p)$ will represent the average consumer valuation for midscale Detroit listings at price $p$. On its part, $\varepsilon$ will represent the value of the “outside” option relative to this segment, accounting for all other Airbnb segments in that city and the option of not purchasing anything within the Airbnb marketplace (e.g., booking an hotel). Our calibration proceeds according to the following steps:
  1. 1.
    We set the total mass of listings $\rho=1$. This allows us to interpret $s^{*}(\lambda,p)$ as the fraction of listings that are available.
  2. 2.
    As usual, we normalize the value of the outside option, $\varepsilon=1$.
  3. 3.
    We set the cost $c=\hat{c}^{(s)}$.
  4. 4.
    We choose a reference price $p$ at which we want to evaluate price experiments. Actual prices were not reported in Farronato and Fradkin, (2022); in lieu of this information, in our results we vary $p$ in a “reasonable” range, as we describe below.
  5. 5.
    We choose an arrival rate $\lambda$. Again, no proxy for $\lambda$ is reported in Farronato and Fradkin, (2022), and so in our results we also vary $\lambda$ in a “reasonable” range, again as we describe below.
  6. 6.
    Finally, given the chosen values of $\rho,\ \varepsilon,\ p$ and $\lambda$, we choose $V$ so that the demand elasticity $pD^{\prime}(p)/D(p)$ for $D(p)=\rho-s^{*}(\lambda,p)$, with $s^{*}(\lambda,p)$ defined in (28), is equal to $\hat{e}_{p}^{(s)}$ (the estimated elasticity from Farronato and Fradkin, (2022)). It is straightforward to check the price elasticity of demand in our demand model is monotonically increasing in $V$, which guarantees that the fitted value of $V$ is uniquely determined in this approach.
Observe that at the conclusion of this procedure, for each value of $\lambda$ and $p$, we have an imputed demand model. We then use this demand model to determine whether a local CR or LR price experiment yields the wrong sign on the estimate of $GTE_{\pi}$, as discussed throughout the paper.
The two “free parameters” in our calibration are the reference price $p$, and the arrival rate $\lambda$, so we vary them in our subsequent empirical analysis. We vary the reference price $p$ in a plausible range from the cost $c$ to a value above the monopoly optimal price implied by $\hat{c}^{(s)}$ and $\hat{e}_{p}^{(s)}$. In our results we also vary $\lambda$ in a plausible range; in particular, we vary $\lambda$ so that in our fitted model we obtain plausible occupancy rates. The occupancy rate is the fraction of total supply that is booked in equilibrium, defined as $\zeta=1-s^{*}(\lambda,p)/\rho$. We vary $\lambda$ to ensure that we plot occupancy rates from 5% to 95%. As an example, one estimate from external market research on short term rental properties Airdna, (2024) estimates that the average occupancy rate across the entire Airbnb market is approximately 60%, with rates closer to 65% for major cities. This suggests the range of values of $\zeta$ in our results includes realistic values.
In our results, we present our visualizations with the two axes of price $p$ and occupancy rate $\zeta$, the latter in place of the arrival rate $\lambda$. We numerically verified that in this model, each pair of original parameters ($\lambda,p$), yields a unique occupancy pair ($\zeta,p$). Further, we have verified that in all instances of our model, as $\lambda$ increases, the resulting $\zeta$ increases monotonically as well. We make this transformation because the occupancy rate is a more natural metric for interpretation than the arrival rate of customers.
8.2 Results
Figure 5: Illustrations of the change of sign region (purple) in LR-experiments (top row) and CR-experiments (bottom row), respectively, for economy listings (left column) and luxury listings (right column) in Boston. The vertical axis is the imputed occupancy rate, and the horizontal axis is the (baseline) control price around which an experiment is conducted. The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red. The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in blue. We plot prices from the cost estimate $\hat{c}^{(s)}$, to a fixed amount above the critical price $p_{c}^{(s)}$.
Figure 5 shows the results of our model’s estimates for LR (top row) and CR (bottom row) price experiments in Boston. The left two plots show results for the economy tier, while the right two plots show results for the luxury tier. We have chosen to include Boston as an example city, but it is important to note that the trends observed in Figure 5, hold for all segments. (See Appendix 11 for these additional figures for all other segments.)
As in our preceding visualizations, the shaded blue region in each plot is where Condition (a) holds in the fitted demand model, cf. (13), and the shaded red region is where Condition (b) holds, cf. (14). These regions overlap in the purple region where both conditions hold. These are the regions where the estimated treatment effect would have the wrong sign. As noted in the preceding section, our results are plotted with the reference price $p$ on the x-axis (varied in a reasonable range, as described in the previous subsection), and the imputed occupancy rate $\zeta$ on the y-axis.
First, observe that in each figure, the right boundary of the change of sign region is a vertical line at a critical price $p_{c}^{(s)}$. To see why, observe that Condition (a) in (13) holds with equality at price $p_{c}^{(s)}$, with the cost $\hat{c}^{(s)}$ and elasticity $\hat{e}_{p}^{(s)}$ corresponding to the segment $s$. In particular, $p_{c}^{(s)}$ satisfies the first order condition for optimal monopoly pricing given $\hat{c}^{(s)}$ and $\hat{e}_{p}^{(s)}$. Further, this identity is insensitive to the value of the occupancy rate $\zeta$ (or equivalently, the arrival rate $\lambda$). Given the monotonicity in Condition (a), at any higher (resp., lower) values of $p$ than $p_{c}^{(s)}$, Condition (a) is guaranteed to not hold (resp., is guaranteed to hold); and again, these facts are also insensitive to the value of $\zeta$.
More importantly, observe that the figures make clear that changes of sign are prevalent in this model. Notably, across the parameter values we tested, LR experiments in particular seem especially vulnerable to changes of sign. This is consistent with the fact that the observed occupancy rates suggest a relatively demand-constrained market, at least as far as bias in price experiments is concerned. On the other hand, CR experiments exhibit a more nuanced behavior. For such experiments, a change of sign is observed in markets that are highly supply-constrained (e.g., with greater than 80-85% of listings being occupied), and when the reference price is near but still below the critical price $p_{c}^{(s)}$.
It should be noted that although CR price experiments exhibit more favorable behavior in these plots than LR price experiments, LR price experiments are more common in industry. In part this is because platforms are often reticent to offer individual customers different prices for the same product (Phillips, 2021). As a consequence, one potential practical takeaway from our work is that firms may in fact have chosen an experimental design that is more commonly producing the wrong sign in their estimated treatment effects in practice, at least in the settings we have studied. Despite the stylized nature of the model, the prevalence of changes of sign across a wide range of parameters suggests this is a phenomenon of significant consequence for practitioners.
9 Conclusion and Future Work
In this paper we studied the biases that arise due to interference in platform experiments, with a particular emphasis on profit and revenue metrics. Notably, we find that in certain settings, these biases can be so severe that they introduce a “change of sign” into the canonical estimators for profit, suggesting to firms that they should move prices (or fees) in the wrong direction. We then derive a novel method for debiasing canonical A/B experiments, by splitting units equally between treatment and control. We conclude by giving empirical evidence for the plausibility of estimators with the wrong sign in practical settings, arguing that firms should be more cautious when using canonical A/B pricing experiments.
In future work we aim to extend our debiasing result, to allow for the correction of canonical experiments in cases where the treatment/control split is not 50/50. In addition to being more flexible, this may allow us to mitigate the sampling inefficiency of our current debiasing method. Currently, our method only uses half of the experimental data, namely, only the observed control demand, which may be a source of increased variance. Finally, we aim to run empirical studies to get a better sense of the true prevalence of changes of sign in practice.